### v2

# Entropy-based random models for hypergraphs

Abstract
Network theory has primarily focused on pairwise relationships, disregarding many-body interactions: neglecting them, however, can lead to misleading representations of complex systems. Hypergraphs represent an increasingly popular alternative for describing polyadic interactions: our innovation lies in leveraging the representation of hypergraphs based on the incidence matrix for extending the entropy-based framework to higher-order structures. In analogy with the Exponential Random Graphs, we name the members of this novel class of models Exponential Random Hypergraphs. Here, we focus on two explicit examples, i.e. the generalisations of the Erdös-Rényi Model and of the Configuration Model. After discussing their asymptotic properties, we employ them to analyse real-world configurations: more specifically, i) we extend the definition of several network quantities to hypergraphs, ii) compute their expected value under each null model and iii) compare it with the empirical one, in order to detect deviations from random behaviours. Differently from currently available techniques, ours is analytically tractable, scalable and effective in singling out the structural patterns of real-world hypergraphs differing significantly from those emerging as a consequence of simpler, structural constraints.
pacs:
89.75.Fb; 02.50.Tt
I I. Introduction
Networks provide a powerful language to model interacting systems Newman (2010); Caldarelli (2010). Within a network framework, the basic unit of interaction, i.e. the edge, involves two nodes and the complexity of the structure as a whole arises from the combination of these units. Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time Lambiotte et al. (2019). Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them Lambiotte et al. (2019); Battiston et al. (2020); Bick et al. (2023). Examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people, brain activities mediated by groups of neurons. Generally speaking, thus, modelling the joint coordination of multiple entities calls for a framework that generalises the traditional, edge-centered one. An increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs as these versatile, mathematical objects naturally allow nodes to interact in groups Battiston et al. (2021).
Given the recent interest towards these objects, the definition of analytical tools to study them is still in its infancy Chodrow (2020); Nakajima et al. (2021); Musciotto et al. (2021); Barthelemy (2022). This paper represents our contribution to fill the gap: hereby, we extend the class of entropy-based null models Jaynes (1957); Park and Newman (2004) to hypergraphs. These models work by preserving a given set of quantities while randomising everything else, hence destroying all, possible correlations between structural properties except for those that are genuinely embodied by the constraints themselves Squartini and Garlaschelli (2011); Cimini et al. (2018); Squartini et al. (2018). The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by spotting out the patterns that are incompatible with simpler, structural constraints Squartini et al. (2013); Saracco et al. (2017); Becatti et al. (2019); Caldarelli et al. (2020); Neal et al. (2021)) or in presence of partial information (to infer the missing portion of a given configuration Parisi et al. (2020)).
Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix. The (generally, rectangular) table contains information about the connectivity of nodes, i.e. the set of hyperedges they belong to, and the connectivity of hyperedges, i.e. the set of nodes they cluster. We will explicitly derive two members of the novel class of models hereby named Exponential Random Hypergraphs, i.e. the Random Hypergraph Model (RHM, generalising the Erdös-Rényi Model) and the Hypergraph Configuration Model (HCM, generalising the Configuration Model), and provide an analytical characterisation of their behaviour. To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph Nakajima et al. (2021); Barthelemy (2022). As a next step, we will employ the HCM to assess the statistical significance of a number of patterns characterising several, real-word hypergraphs.
II II. Formalism and basic quantities
A hypergraph can be defined as a pair $H(\mathcal{V},\mathcal{E}_{H})$ where $\mathcal{V}$ is the set of vertices and $\mathcal{E}_{H}$ is the set of hyperedges. Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in Berge (1967, 1973), where hyperedges ‘link’ one or more vertices, and in Voloshin (2009), where hyperedges are allowed to be empty sets as well. Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: in words, we consider non-simple hypergraphs, admitting loops and parallel edges (i.e. hyperedges involving exactly the same set nodes) of any size, including 0 (corresponding to empty hyperedges), and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
$\mathbf{I}\>=\>\begin{tabular}[]{c|c|c|c|c|c|}&$\mathbf{e_{1}}$&$\mathbf{e_{2}}$&$\mathbf{e_{3}}$&$\mathbf{e_{4}}$&$\mathbf{e_{5}}$\\ \hline\cr$\mathbf{n_{1}}$&0&1&1&0&1\\ \hline\cr$\mathbf{n_{2}}$&0&0&1&1&1\\ \hline\cr$\mathbf{n_{3}}$&0&1&1&1&0\\ \hline\cr$\mathbf{n_{4}}$&1&0&0&1&1\\ \hline\cr$\mathbf{n_{5}}$&1&1&1&1&0\\ \hline\cr$\mathbf{n_{6}}$&0&1&0&0&0\\ \hline\cr\end{tabular}$
Table 1: Example of incidence matrix describing a binary, undirected hypergraph with $6$ nodes and $5$ hyperedges. The generic entry $I_{i\alpha}$ is 1 if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise (see also Fig. 1).
As for traditional graphs, an algebraic representation of hypergraphs can be devised as well. For a formal analogy with the traditional case, let us call the cardinality of the set of nodes $|\mathcal{V}|=N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|=L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge. If we indicate the incidence matrix with $\mathbf{I}$, its generic entry $I_{i\alpha}$ will be $1$ if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise. Notice that the number of $1$s along each row can vary between 00 and $L$, the former case indicating an isolated node and the latter one indicating a node that belongs to each hyperedge; similarly, the number of $1$s along each column can vary between 00 and $N$, the former case indicating an empty hyperedge and the latter one indicating a hyperedge that includes all nodes. As explicitly noticed elsewhere Nakajima et al. (2021); Barthelemy (2022), representing a hypergraph via its incidence matrix is equivalent at considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$. Table 1 represents the incidence matrix describing the binary, undirected hypergraph shown in Fig. 1.
Once the incidence matrix has been defined, several quantities needed for the description of hypergraphs can be defined quite straightforwardly: for example, the ‘degree of node $i$’ (hereby, degree) reads
$k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}$ (1)
and counts the number of hyperedges that are incident to it; analogously, the ‘degree of hyperedge $\alpha$’ (hereby, hyperdegree) reads
$h_{\alpha}=\sum_{i=1}^{N}I_{i\alpha}$ (2)
and counts the number of nodes it clusters. Both the sum of degrees and that of hyperdegrees equal the total number of $1$s, i.e. $\sum_{i=1}^{N}k_{i}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}=\sum_{\alpha=1}^{L}\sum_{i=1}^{N}I_{i\alpha}=\sum_{\alpha=1}^{L}h_{\alpha}\equiv T$. Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information on the hyperedge size.
Figure 1: Cartoon representation of the hypergraph whose incidence matrix is shown in Table 1.
III III. Binary, undirected hypergraphs randomisation
An early attempt to define randomisation models for hypergraphs can be found in Ghoshal et al. (2009). Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network Nakajima et al. (2021). Considering the incidence matrix has two, clear advantages over the tensor-based representation employed in Ghoshal et al. (2009); Nakajima et al. (2021): i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
In order to extend the rich set of null models induced by graph-specific, global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting. In what follows, we will consider the total number of 1s, i.e. $T$, the degree and the hyperdegree sequences, i.e. $\{k_{i}\}_{i=1}^{N}$ and $\{h_{\alpha}\}_{\alpha=1}^{L}$ either separately or in a joint fashion; moreover, we will distinguish between microcanonical and canonical randomisation techniques.
III.1 A. Homogeneous benchmarks:
the Random Hypergraph Model (RHM)
The model is defined by just one, global constraint, that, in our case, reads $T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}$. Its microcanonical version extends the model by Erdös and Rényi Erdös and Rényi (1959) - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of 1s, say $T^{*}$: they are
$\Omega_{\text{RHM}}=\binom{V}{T^{*}}$ (3)
with $V=NL$ being the total number of entries of the incidence matrix $\mathbf{I}$. Once the total number of configurations composing the microcanonical ensemble is determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of 1s is kept fixed while any other correlation is destroyed.
The canonical version of the RHM, instead, extends the model by Gilbert Gilbert (1959) and rests upon the constrained maximisation of Shannon entropy, i.e.
$\mathscr{L}\equiv S[P]-\sum_{i=0}^{M}\theta_{i}[P(\mathbf{I})C_{i}(\mathbf{I})-\langle C_{i}\rangle]$ (4)
where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties. The sum defining Shannon entropy runs over the set $\mathcal{I}$ of incidence matrices described in the introductory paragraph and known as canonical ensemble.
Such an optimisation procedure defines the Exponential Random Hypergraph (ERH) framework, described by the expression
$P(\mathbf{I})=\frac{e^{-H(\mathbf{I})}}{Z}=\frac{e^{-H(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-H(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}.$ (5)
In the simplest case, the only, global constraint is represented by $T$ and leads to the expression
$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}$
$=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (6)
that can be re-written as
$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}p^{I_{i\alpha}}(1-p)^{1-I_{i\alpha}}=p^{T(\mathbf{I})}(1-p)^{NL-T(\mathbf{I})}$ (7)
with $e^{-\theta}\equiv x$ and $p\equiv x/(1+x)$. The canonical ensemble, now, includes all $N\times L$ rectangular matrices whose number of entries equating 1 ranges from 0 to $NL$. According to such a model, the entries of the incidence matrix are i.i.d. Bernoulli random variables, i.e. $I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$: as a consequence, the total number of 1s, the degrees and the hyperdegrees obey Binomial distributions - being all defined as sums of i.i.d. Bernoulli random variables. Specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
In order to ensure that $\langle T\rangle_{\text{RHM}}=T^{*}$, parameters have to be tuned opportunely. To this aim, the likelihood maximisation principle can be invoked Garlaschelli and Loffredo (2008): it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter(s) that define it. Such a recipe leads us to find $p=T^{*}/NL$, with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
The RHM (also considered in Barthelemy (2022), although a derivation from first principles is missing, there) is formally equivalent to the Bipartite Random Graph Model (BiRGM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.1.1 1. Estimation of the number of empty hyperedges
As non-simple hypergraphs admit the presence of empty as well as parallel hyperedges, let us evaluate how frequently they appear in the ensembles induced by our benchmarks.
To this aim, let us start by considering the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero. Upon remembering that $h_{\alpha}\sim\text{Bin}(N,p)$, one finds
$p_{0}\equiv(1-p)^{N}.$ (8)
Such a quantity leads us to find the probability of observing at least one, empty hyperedge (or, in other terms, the complementary of the probability that no hyperedge is empty), i.e.
$p_{1}\equiv 1-(1-p_{0})^{L}=1-[1-(1-p)^{N}]^{L},$ (9)
and the expected number of empty hyperedges: upon remembering that they are i.i.d. Binomial random variables, one finds
$\langle N_{0}\rangle\equiv Lp_{0}=L(1-p)^{N}.$ (10)
Let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the RHM, as the density of 1s in the incidence matrix, i.e. $p=T/NL$, varies. The two regimes of interest are the dense one, defined by $T\rightarrow NL$, and the sparse one, defined by $T\rightarrow 0$. In the dense case, one finds
$\lim_{T\to NL}p_{0}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (11)
a relationship inducing $p_{1}\xrightarrow[]{T\to NL}0$ and $\langle N_{0}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{T\to 0}p_{0}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (12)
a relationship inducing $p_{1}\xrightarrow[]{T\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
To evaluate the density of 1s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, the probability for a generic hyperedge to be empty remains finite as well: consistently, the expected number of empty hyperedges becomes
$\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}L\left(1-\frac{h}{N}\right)^{N}=Le^{-h};$ (13)
upon imposing $Le^{-h}\leq 1$, i.e. that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}=\ln L$. In words, a value $p>p_{f}^{\text{RHM}}\equiv h_{f}^{\text{RHM}}/N=\ln L/N$ ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one. The calculation above complements the result derived in Barthelemy (2022) about the percolation threshold for hypergraphs (see also Appendix A).
III.1.2 2. Estimation of the number of parallel hyperedges
Let us, now, move to considering the issue of parallel hyperedges. By definition, two, parallel hyperedges $\alpha$ and $\beta$ are characterised by identical columns: hence, their Hamming distance, defined as the number of positions at which the corresponding symbols are different, is zero. More formally,
$d_{\alpha\beta}\equiv\sum_{i=1}^{N}[I_{i\alpha}(1-I_{i\beta})+I_{i\beta}(1-I_{i\alpha})],$ (14)
a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$. Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that $P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ and $\langle d_{\alpha\beta}\rangle=2p(1-p)N$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{T\to NL}P(d_{\alpha\beta}=0)$ $=\lim_{T\to 0}P(d_{\alpha\beta}=0)=1$ (15)
and
$\lim_{T\to NL}\langle d_{\alpha\beta}\rangle$ $=\lim_{T\to 0}\langle d_{\alpha\beta}\rangle=0$ (16)
as $(2T/NL)(1-T/NL)\to 0$ in both regimes: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
Let us, now, evaluate the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$, within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e. $p=O(1/N)$. Since $p=T/NL$ and $h\equiv T/L$ remains finite, one finds that
$\lim_{N\to+\infty}\langle d_{\alpha\beta}\rangle=\lim_{N\to+\infty}\frac{2h}{N}\left(1-\frac{h}{N}\right)N=2h;$ (17)
upon imposing $2h\geq 1$, i.e. that the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}=1/2$. In words, a value $p>p_{r}^{\text{RHM}}\equiv h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element (see also Appendix A).
III.1.3 3. Estimation of the overlap between hyperedges
Finally, let us move to considering the issue of overlapping hyperedges. To this aim, let us require that the probability for any, two hyperedges to overlap amounts to 1. In formulas,
$1-(1-p^{2})^{N}=1,$ (18)
a relationship leading to the equation
$(1-p^{2})^{N}\simeq 1-Np^{2}=0$ (19)
that holds true in the sparse regime and further implies
$h^{2}=N;$ (20)
such a value, firstly derived in Barthelemy (2022), individuates the so-called percolation threshold, corresponding to $h_{p}^{\text{RHM}}=\sqrt{N}$. In words, a value $p>p_{p}^{\text{RHM}}\equiv h_{p}^{\text{RHM}}/N=1/\sqrt{N}$ ensures that any, two hyperedges in our random hypergraph share at least one node.
Since $h_{r}^{\text{RHM}}\leq h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, such a result contributes to describe the following situation: by progressively rising the parameter $p$, three, different thresholds are met, i.e. i) the resolution threshold $p_{r}^{\text{RHM}}=1/2N$, ensuring that, on average, any, two hyperedges differ by at least one element; ii) the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, ensuring that the expected number of empty hyperedges is strictly less than one; iii) the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{N}$, ensuring that any, two hyperedges overlap.
III.2 B. Heterogeneous benchmarks:
the Hypergraph Configuration Model (HCM)
The number of constraints can be enlarged to include the degrees, i.e. the sequence $\{k_{i}\}_{i=1}^{N}$, and the hyperdegrees, i.e. the sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$. Counting the number of configurations on which both sequences match their empirical values is a hard task - although numerical recipes to shuffle the entries of a rectangular matrix, while preserving its marginals, exist Strona et al. (2014); Carstens (2015); Chodrow (2020). Solving the corresponding problem in the canonical framework is, instead, straightforward. Indeed, Shannon entropy maximisation leads to
$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}$
$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
$=\prod_{i=1}^{N}x_{i}^{k_{i}(\mathbf{I})}\prod_{\alpha=1}^{L}y_{\alpha}^{h_{\alpha}(\mathbf{I})}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x_{i}y_{\alpha})^{-1},$ (21)
an expression that can be re-written as
$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}p_{i\alpha}^{I_{i\alpha}}(1-p_{i\alpha})^{1-I_{i\alpha}}$ (22)
with $e^{-\alpha_{i}}\equiv x_{i}$, $\forall\>i$, $e^{-\beta_{\alpha}}\equiv y_{\alpha}$, $\forall\>\alpha$ and $p_{i\alpha}\equiv x_{i}y_{\alpha}/(1+x_{i}y_{\alpha})$, $\forall\>i,\alpha$. According to such a model, the entries of the incidence matrix of a hypergraph are independent random variables that obey different Bernoulli distributions, i.e. $I_{i\alpha}\sim\text{Ber}(p_{i\alpha})$, $\forall\>i,\alpha$. As a consequence, both degrees and hyperdegrees obey Poisson-Binomial distributions, i.e. $k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ Saracco et al. (2017).
In this case, solving the likelihood maximisation problem amounts to solve the system of coupled equations
$k_{i}^{*}$ $=\sum_{\alpha=1}^{L}\frac{x_{i}y_{\alpha}}{1+x_{i}y_{\alpha}}=\sum_{\alpha=1}^{L}p_{i\alpha}=\langle k_{i}\rangle,\>\forall\>i$ (23)
$h_{\alpha}^{*}$ $=\sum_{i=1}^{N}\frac{x_{i}y_{\alpha}}{1+x_{i}y_{\alpha}}=\sum_{i=1}^{N}p_{i\alpha}=\langle h_{\alpha}\rangle,\>\forall\>\alpha$ (24)
ensuring that $\langle k_{i}\rangle=k_{i}^{*}$, $\forall\>i$, $\langle h_{\alpha}\rangle=h_{\alpha}^{*}$, $\forall\>\alpha$ (and, as a consequence, $\langle T\rangle=T^{*}$). In case hypergraphs are sparse and in absence of hubs
$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (25)
The HCM reduces to a ‘partial’ Configuration Model Saracco et al. (2017) when either the degree or the hyperdegree sequence is left unconstrained (see also Appendix B). The canonical ensemble of each randomisation model (Table 2 in Appendix B sums up the set of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
III.2.1 1. Estimation of the number of empty hyperedges
Let us, now, consider the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero. Upon remembering that $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, one finds
$p_{0}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha}).$ (26)
Such a quantity leads to the probability of observing at least one, empty hyperedge, now reading
$p_{1}\equiv 1-\prod_{\alpha=1}^{L}\left[1-p_{0}^{\alpha}\right]=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right],$ (27)
and to the expected number of empty hyperedges
$\langle N_{0}\rangle\equiv\sum_{\alpha=1}^{L}p_{0}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (28)
As in the previous section, let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM, as the density of 1s in the incidence matrix varies. Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can still be drawn within a simpler framework. To this aim, let us consider the functional form reading
$p_{i\alpha}=\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}},\>\forall\>i,\alpha$ (29)
where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of 1s in the incidence matrix [^1] Partial Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$.. Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution. The dense and sparse regimes are, now, defined by the positions $z\to+\infty$ and $z\to 0$, respectively. In the dense case, one finds
$\lim_{z\to+\infty}p_{0}^{\alpha}$ $=\lim_{z\to+\infty}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (30)
a relationship inducing $p_{1}\xrightarrow[]{z\to+\infty}0$ and $\langle N_{0}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases. Consistently, in the sparse case one finds
$\lim_{z\to 0}p_{0}^{\alpha}$ $=\lim_{z\to 0}\prod_{i=1}^{N}(1-p_{i\alpha})$
$=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (31)
a relationship inducing $p_{1}\xrightarrow[]{z\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{z\to 0}L$. In words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
III.2.2 2. Estimation of the number of parallel hyperedges
Let us, now, focus on the issue of parallel hyperedges. As in the case of the RHM, we consider the Hamming distance between the columns representing the two hyperedges $\alpha$ and $\beta$. Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ and $\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$, $\forall\>\alpha\neq\beta$. Notice that
$\lim_{z\to+\infty}P(d_{\alpha\beta}=0)$ $=\lim_{z\to 0}P(d_{\alpha\beta}=0)=1$ (32)
and
$\lim_{z\to 0}\langle d_{\alpha\beta}\rangle$ $=\lim_{z\to 0}\langle d_{\alpha\beta}\rangle=0$ (33)
as $q_{i}^{\alpha\beta}\to 0$ in both regimes. As in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s. Analogously for the expected Hamming distance.
When considering heterogeneous benchmarks, closed-form expressions for the asymptotic estimates can be provided only in specific cases (see also Appendix B).
Figure 2: Left panel: trend of $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e. the probability of observing at least one empty hyperedge as a function of $T/NL$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (whose related, critical density value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ individuates a steep transition separating the sparse regime from the dense one) returns the value $0.63231-(1-1/L)^{L}=$ 0.6323. Right panel: trend of $p_{1}$ as defined by Eq. (27). Notice that the trend induced by the HCM is shifted on the right with respect to the trend induced by the RHM: this is probably due to the presence of small fitnesses increasing the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
IV IV. Results
IV.1 A. Hypergraphs in the dense and sparse regime
Let us start by verifying the correctness of the estimations of the number of empty and parallel hyperedges provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$ (for other values, see Appendix A and Appendix B).
IV.1.1 1. The Random Hypergraph Model
Let us start focusing on our homogeneous benchmark. The left panel of Fig. 2 depicts the trend of $p_{1}=1-\left[1-\left(1-\frac{T}{NL}\right)^{N}\right]^{L}$. The left panel of Fig. 3 depicts the trend of $\frac{\langle N_{0}\rangle}{L}=\left(1-\frac{T}{NL}\right)^{N}=p_{0}$ and $P(d_{\alpha\beta}=0)=\left[1-\frac{2T}{NL}\left(1-\frac{T}{NL}\right)\right]^{N}$. Each quantity has been plotted as a function of $T/NL\in[10^{-6},1]$. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval. A first, general comment concerns the agreement between the analytical trends (solid lines) and the numerical estimations (dots), confirming the correctness of our formulas.
The filling threshold.
Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{0}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1-(1-1/L)^{L}\simeq 0.6323$ interestingly enough, letting $L$ grow asymptotically as well, the previous value tends to $1-e^{-1}\simeq 0.6321$.
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $0.0011/L=$ 0.001: in words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e. 2/3$\simeq$ 2/3)≃ 2 / 3 ). The steepness of the trend of $p_{1}$, however, suggests it to quickly vanish as the density of 1s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
Figure 3: Left panel: trends of $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$ (aquamarine line) and $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ (dark-blue line) as a function of $T/NL$; evaluating them in correspondence of $h_{r}^{\text{RHM}}=1/2$ (the related, critical density value is indicated in light-blu) returns the values $e^{-{h_{r}^{\text{RHM}}}}=1/\sqrt{e}\simeq 0.606$ and $e^{-{2h_{r}^{\text{RHM}}}}=e^{-1}\simeq 0.367$; evaluating them in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (the related, critical density value is indicated in magenta) returns the values $e^{-{h_{f}^{\text{RHM}}}}=1/L=0.001$ and $e^{-{2h_{f}^{\text{RHM}}}}=1/L^{2}=10^{-6}$. The red line indicates the density value corresponding to $h=1$, ensuring that each, expected hyperdegree amounts to 1. Right panel: trends of $p_{0}$, as defined by Eq. (26), and $P_{0}$, i.e. the percentage of pairs of parallel hyperedges. As already noticed, the shift on the right of the trend induced by the HCM is probably due to the presence of small fitnesses, decreasing the magnitude of the set of probability coefficients. The dense (sparse) regime is recovered for large (small) values of $T/NL$. Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval. The yellow line indicates the density value in correspondence of which the minimum, expected hyperdegree amounts to 1.
The resolution threshold.
The value of the resolution threshold has been determined by inspecting the asymptotic behaviour of $\langle d_{\alpha\beta}\rangle$: evaluating $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $e^{-1}\simeq 0.367$.
Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $1/\sqrt{e}\simeq 0.606$: in words, the resolution threshold ensures the likelihood of observing any, two parallel hyperedges to be almost half of the probability of each, single hyperedge to be empty, hence suggesting that getting rid of parallel hyperedges is ‘easier’ than getting rid of empty ones.
Crossing the thresholds.
As Fig. 3 confirms, the probability for any, two hyperedges to be parallel decreases faster than the probability of each, single hyperedge to be empty: in words, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges, when empty hyperedges are no longer observed, parallel hyperedges are no longer observed, too. More formally, $h_{r}^{\text{RHM}}=1/2<h_{f}^{\text{RHM}}=\ln L$: hence, progressively rising the density of 1s in the incidence matrix leads us crossing the resolution threshold before crossing the filling threshold. In more quantitative terms, evaluating $\langle N_{0}\rangle\xrightarrow[]{N\to+\infty}Le^{-h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $L/\sqrt{e}\simeq 606$ which is larger than $\langle d_{\alpha\beta}\rangle\xrightarrow[]{N\to+\infty}2h$ evaluated in correspondence of $h_{r}^{\text{RHM}}=1/2$ returning the value 1 by definition.
IV.1.2 2. The Hypergraph Configuration Model
Let us, now, focus on the trends induced by our heterogeneous benchmarks. Following the procedure described in the previous sections, both the fitnesses of nodes and those of hyperedges were drawn from a Pareto distribution [^2] Other fat-tailed distributions were considered: qualitatively, results do not change. with $\alpha=2$. The right panel of Fig. 2 depicts the trend of $p_{1}$, as defined by Eq. (27). The right panel of Fig. 3 depicts the trend of $p_{0}$, as defined by Eq. (26), and the arithmetic mean $P_{0}\equiv 2\sum_{\alpha<\beta}P(d_{\alpha\beta}=0)/L(L-1)$ with $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$, i.e. the average of the value $P(d_{\alpha\beta}=0)$ over all pairs of hyperedges - a choice dictated by the will of accounting for the heterogeneity of different pairs of hyperedges. Both quantities have been plotted as a function of $\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$. The dense (sparse) regime is recovered for large (small) values of $z$. Each dot of Figs. 2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval. As in the case of the RHM, the agreement between the analytical trends (solid lines) and the numerical estimations (dots) confirms the correctness of our formulas.
Comparing the analytical estimation of (the trend of) the percentage of pairs of parallel hyperedges with its sample counterpart reveals the former to provide a quite accurate estimation of the latter. Besides, we explicitly notice that the trends characterising the HCM are sistematically shifted on the right: this is probably due to the presence of small fitnesses that, at least in case of Fig. 2, increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish. Such a result seems to be confirmed by the trends depicted in Fig. 3.
Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of 1s in the incidence matrix guaranteeing that the minimum, expected hyperdegree amounts to 1: as evident from Fig. 3, such a threshold still lies in the right tail of the trends induced by the HCM - although its, precise, numerical value depends on those of the fitnesses.
IV.2 B. Solving the HCM on real-world hypergraphs
In order to test our benchmarks on real-world configurations, we have focused on a number of data sets taken from Austin R. Benson’s website [^3] https://www.cs.cornell.edu/~arb/data/, i.e. the contact-primary-school, the email-Enron and the NDC-classes ones.
To have an explicit expression for the expected value of any quantity of interest, let us comment on the goodness of the approximation
$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (34)
As Fig. 7 in Appendix A shows, it is is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
IV.3 C. ‘Hypergraph to graph’ projection
The canonical formalism that we have adopted leads to factorisable distributions, i.e. distributions that can be written as a product of pair-wise probability distributions; this allows the expectation of several quantities of interest to be evaluated analytically. To this aim, let us consider the matrix introduced in Battiston et al. (2020), reading
$\mathbf{W}=\mathbf{I}\cdot\mathbf{I}^{T}-\mathbf{K}$ (35)
with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above
$w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}-\delta_{ij}k_{i}$ (36)
counts the number of hyperedges nodes $i$ and $j$ are connected by - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$. The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
$\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}\rangle\langle I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha};$ (37)
the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 4) can be computed as
$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (38)
whose expected value reads
$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (39)
As further confirmed by Fig. 8 in Appendix A, the approximation provided by Eq. (34) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$. Interestingly, as Fig. 5a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
Let us, now, extend the concept of assortativity to hypergraphs. To this aim, we consider the quantity named average incident hyperedges degree, defined as
$k_{i}^{nn}=\sum_{\alpha=1}^{L}\frac{I_{i\alpha}h_{\alpha}}{k_{i}}=\frac{\sigma_{i}+k_{i}}{k_{i}}=\frac{\sigma_{i}}{k_{i}}+1\simeq\frac{\sigma_{i}}{k_{i}}$ (40)
and representing the arithmetic mean of the degrees of the hyperedges including node $i$. An analytical approximation of its expected value can be provided as well:
$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}$
$=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (41)
Figure 4: Graphical representation of the ‘hypergraph to graph’ projection. Notice that while the degree of each node (i.e. the number of hyperedges that are incident to it) reads $\{k_{i}\}=(3,3,3,3,4,1)$, the sigma of each node (i.e. the total number of hyperedges shared by it with any other node or, equivalently, the strength on the projection) reads $\{\sigma_{i}\}=(8,8,9,6,10,3)$ and the kappa of each node (i.e. the number of nodes it shares an hyperedge with or, equivalently, the degree on the projection) reads $\{\kappa_{i}\}=(5,4,5,4,5,3)$.
IV.4 D. Disparity ratio and degree in the projection
More information about the patterns shaping real-world hypergraphs can be obtained upon defining the ratio $f_{ij}=w_{ij}/\sigma_{i}$, $i\neq j$ that induces the quantity
$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (42)
known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig. 4 and Fig. 5b). Since, under the RHM, $w_{ij}\sim\text{Bin}(L,p^{2})$, we find that $\langle a_{ij}\rangle=1-(1-p^{2})^{L}$, i.e. the expected value of $a_{ij}$ coincides with the probability of observing a non-zero overlap. Under the HCM, instead, $w_{ij}\sim\text{PoissBin}(L,\{p_{i\alpha}p_{j\alpha}\}_{\alpha=1}^{L})$, hence
$\langle a_{ij}\rangle=1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha}).$ (43)
Let us also notice that
$Y_{i}=\frac{1}{\kappa_{i}}$ (44)
in case weights are equally distributed among the connections established by node $i$, i.e. $w_{ij}=a_{ij}\sigma_{i}/\kappa_{i}$, $i\neq j$. Any larger value signals an excess concentration of weight in one or more links. An analytical approximation of the expected value of the disparity ratio of node $i$ can be provided as well:
$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (45)
Figure 5: Scatter plots between the empirical values $\{\sigma_{i}\}$, $\{\kappa_{i}\}$, $\{Y_{i}\}$, $\{\text{CEC}_{i}\}$ and the expected values $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$, $\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$, $\langle\{Y_{i}\rangle_{\text{HCM}}\}$, $\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$. The HCM overestimates the extent to which any, two nodes overlap, as well as the CEC; the disparity ratio, instead, is underestimated by it. These results can be understood by considering that the HCM just constrains the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed.
Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed. In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig. 10 in Appendix C). For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM. In any case, as Fig. 5c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected. A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})]$: as Fig. 5b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected. This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs LaRock and Lambiotte (2023); Landry et al. (2024).
IV.5 F. Eigenvector centrality
Centrality measures for hypergraphs have been defined as well. An example is provided by the clique motif eigenvector centrality (CEC), defined in Benson (2019) (see also Appendix D): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$. As Fig. 5d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
Figure 6: Top panels: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes. Bottom panels: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges. Communities have been detected by running the Louvain algorithm.
IV.6 G. Confusion matrix
Let us, now, consider the set of properties composing the so-called confusion matrix (see also Appendix E). They are known as the true positive rate (TPR), i.e. the percentage of 1s correctly recovered by a given method and whose expected value reads
$\langle\text{TPR}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{T};$ (46)
the specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method and whose expected value reads
$\langle\text{SPC}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{(1-I_{i\alpha})(1-p_{i\alpha})}{NL-T};$ (47)
the positive predictive value (PPV), i.e. the percentage of 1s correctly recovered by a given method with respect to the total number of 1s predicted by it and whose expected value reads
$\langle\text{PPV}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{\langle T\rangle};$ (48)
and the accuracy (ACC), measuring the overall performance of a given method in correctly placing both 1s and 0s and whose expected value reads
$\langle\text{ACC}\rangle=\frac{\langle\text{TP}\rangle+\langle\text{TN}\rangle}{NL}.$ (49)
Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 3 in Appendix E); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of 1s - and, as a consequence, the density of 0s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 4 in Appendix F).
IV.7 H. Community detection
Communities are commonly understood as densely connected groups of nodes. Representing an hypergraph via its incidence matrix allows this statement to be made more precise from a statistical perspective: in fact, the null models discussed so far can be employed to test if any two nodes share a significantly large number of hyperedges - hence can be clustered together, should this be the case.
To this aim, we can adapt the recipe proposed in Saracco et al. (2017) to project bipartite networks and summarised in the following. One, first, computes
[…]
where $n=N(N-1)/2$ and $t$ is the single-test significance level, set to $0.01$ in the present analysis. Third, one links the (pairs of) nodes whose related p-value is smaller than the aforementioned threshold.
Figure 6 shows the result returned by the Louvain algorithm, run on the validated projections: as noticed elsewhere Saracco et al. (2017); Pratelli et al. (2024); Guarino et al. (2024), the detection of mesoscale structures is enhanced if carried out on filtered topologies.
V V. Conclusions
Our paper contributes to current research on hypergraphs by extending the constrained entropy-maximisation framework to incidence matrices, i.e. their simplest, tabular representation. Differently from the currently-available techniques Chodrow (2020), our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
Beside leading to results whose relevance is mostly theoretical (i.e. the individuation of different regimes for higher-order structures and the estimation of the actual impact of empty and parallel hyperedges on the analysis of empirical systems), our models prove to be particularly useful when employed as benchmarks for real-world systems, i.e. for detecting patterns that are not imputable to purely random effects. Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Appendix F).
This is even more surprising when considering that our results are obtained under a benchmark such as the HCM, i.e. a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes de Arruda et al. (2021); Iacopini et al. (2022); St-Onge et al. (2022) - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
VI VI. Acknowledgements
RL acknowledges support from the EPSRC grants EP/V013068/1, EP/V03474X/1 and EP/Y028872/1. TS acknowledges support from SoBigData.it that receives funding from European Union – NextGenerationEU – National Recovery and Resilience Plan (Piano Nazionale di Ripresa e Resilienza, PNRR) – Project: ‘SoBigData.it – Strengthening the Italian RI for Social Mining and Big Data Analytics’ - Prot. IR0000013 – Avviso n. 3264 del 28/12/2021. FS acknowledges support from the project ‘CODE - Coupling Opinion Dynamics with Epidemics’, funded under PNRR Mission 4 ‘Education and Research’ - Component C2 - Investment 1.1 - Next Generation EU ‘Fund for National Research Program and Projects of Significant National Interest’ PRIN 2022 PNRR, grant code P2022AKRZ9.

### v3

@@
-Entropy-based random models for hypergraphs
+
 Abstract
-Network theory has primarily focused on pairwise relationships, disregarding many-body interactions: neglecting them, however, can lead to misleading representations of complex systems.
-Hypergraphs represent an increasingly popular alternative for describing polyadic interactions: our innovation lies in leveraging the representation of hypergraphs based on the incidence matrix for extending the entropy-based framework to higher-order structures.
-In analogy with the Exponential Random Graphs, we name the members of this novel class of models Exponential Random Hypergraphs.
-Here, we focus on two explicit examples, i.e.
-the generalisations of the Erdös-Rényi Model and of the Configuration Model.
-After discussing their asymptotic properties, we employ them to analyse real-world configurations: more specifically, i) we extend the definition of several network quantities to hypergraphs, ii) compute their expected value under each null model and iii) compare it with the empirical one, in order to detect deviations from random behaviours.
-Differently from currently available techniques, ours is analytically tractable, scalable and effective in singling out the structural patterns of real-world hypergraphs differing significantly from those emerging as a consequence of simpler, structural constraints.
+Network theory has often disregarded many-body relationships, solely focusing on pairwise interactions: neglecting them, however, can lead to misleading representations of complex systems.
+Hypergraphs represent a suitable framework for describing polyadic interactions.
+Here, we leverage the representation of hypergraphs based on the incidence matrix for extending the entropy-based approach to higher-order structures: in analogy with the Exponential Random Graphs, we introduce the Exponential Random Hypergraphs (ERHs).
+After exploring the asymptotic behaviour of thresholds generalising the percolation one, we apply ERHs to study real-world data.
+First, we generalise key network metrics to hypergraphs; then, we compute their expected value and compare it with the empirical one, in order to detect deviations from random behaviours.
+Our method is analytically tractable, scalable and capable of revealing structural patterns of real-world hypergraphs that differ significantly from those emerging as a consequence of simpler constraints.
+keywords:
+complex networks, higher-order structures, exponential random hypergraphs
 pacs:
 89.75.Fb; 02.50.Tt
-I I. Introduction
-Networks provide a powerful language to model interacting systems Newman (2010); Caldarelli (2010).
-Within a network framework, the basic unit of interaction, i.e.
+Introduction
+Networks provide a powerful language to model interacting systems [1, 2].
+Within such a framework, the basic unit of interaction, i.e.
 the edge, involves two nodes and the complexity of the structure as a whole arises from the combination of these units.
-Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time Lambiotte et al.
-(2019).
-Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them Lambiotte et al.
-(2019); Battiston et al.
-(2020); Bick et al.
-(2023).
-Examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people, brain activities mediated by groups of neurons.
-Generally speaking, thus, modelling the joint coordination of multiple entities calls for a framework that generalises the traditional, edge-centered one.
-An increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs as these versatile, mathematical objects naturally allow nodes to interact in groups Battiston et al.
-(2021).
-Given the recent interest towards these objects, the definition of analytical tools to study them is still in its infancy Chodrow (2020); Nakajima et al.
-(2021); Musciotto et al.
-(2021); Barthelemy (2022).
-This paper represents our contribution to fill the gap: hereby, we extend the class of entropy-based null models Jaynes (1957); Park and Newman (2004) to hypergraphs.
-These models work by preserving a given set of quantities while randomising everything else, hence destroying all, possible correlations between structural properties except for those that are genuinely embodied by the constraints themselves Squartini and Garlaschelli (2011); Cimini et al.
-(2018); Squartini et al.
-(2018).
-The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by spotting out the patterns that are incompatible with simpler, structural constraints Squartini et al.
-(2013); Saracco et al.
-(2017); Becatti et al.
-(2019); Caldarelli et al.
-(2020); Neal et al.
-(2021)) or in presence of partial information (to infer the missing portion of a given configuration Parisi et al.
-(2020)).
-Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix.
-The (generally, rectangular) table contains information about the connectivity of nodes, i.e.
-the set of hyperedges they belong to, and the connectivity of hyperedges, i.e.
-the set of nodes they cluster.
-We will explicitly derive two members of the novel class of models hereby named Exponential Random Hypergraphs, i.e.
+Despite its many successes, network science disregards certain aspects of interacting systems, notably the possibility that more-than-two constituent units could interact at a time [3].
+Yet, it has been increasingly shown that, for a variety of systems, interactions cannot be always decomposed into a pairwise fashion and that neglecting higher-order ones can lead to an incomplete, if not misleading, representation of them [3, 4, 5]: examples include chemical reactions involving several compounds, coordination activities within small teams of co-working people and brain activities mediated by groups of neurons.
+Generally speaking, thus, modelling the joint coordination of multiple entities calls for a generalisation of the traditional edge-centred framework.
+While approaches focusing on the so-called simplicial complexes have been proposed [6, 7], an increasingly popular alternative to support a science of many-body interactions is provided by hypergraphs, as these mathematical objects allow nodes to interact in groups without posing restrictions such as the ‘hierarchical’ ones characterising the former ones [8] - which, in fact, include all the subsets of a given simplex [6].
+Several contributions to the definition of analytical tools for their study have already appeared [9, 10, 11, 12]: while some pertain to the purely mathematical literature and have considered probabilistic hypergraphs with the aim of studying properties such as the existence of cycles, cliques, etc.
+[13, 14, 15], others have adopted approaches rooted into statistical physics.
+Among the latter ones, some have proposed microcanonical approaches [9, 11] while others have focused on canonical ones [10, 16, 12].
+The present contribution aims at extending the class of entropy-based benchmarks [17, 18, 19, 20] to hypergraphs while providing a coherent framework to formally derive the canonical approaches that have been proposed so far.
+These models work by preserving a given set of quantities while randomising everything else, hence destroying all possible correlations between structural properties except for those that are genuinely embodied into the constraints themselves [20, 21, 22].
+The versatility of such an approach allows it to be employed either in presence of full information (to quantify the level of self-organisation of a given configuration by identifying the patterns that are incompatible with simpler, structural constraints [23, 24, 25, 26, 27]) or in presence of partial information (to infer the missing portion of a given configuration [28]).
+Our strategy for defining null models for hypergraphs is based on the randomisation of their incidence matrix, i.e.
+the (generally, rectangular) table contains information about the connectivity of nodes (the set of hyperedges they belong to) and the connectivity of hyperedges (the set of nodes they cluster).
+We will explicitly derive two members of this novel class of models hereby named Exponential Random Hypergraphs, i.e.
 the Random Hypergraph Model (RHM, generalising the Erdös-Rényi Model) and the Hypergraph Configuration Model (HCM, generalising the Configuration Model), and provide an analytical characterisation of their behaviour.
-To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph Nakajima et al.
-(2021); Barthelemy (2022).
-As a next step, we will employ the HCM to assess the statistical significance of a number of patterns characterising several, real-word hypergraphs.
-II II. Formalism and basic quantities
+To this aim, we will exploit the formal equivalence between the incidence matrix of a hypergraph and the biadjacency matrix of a bipartite graph [10, 12].
+Afterwards, we will employ the HCM to assess the statistical significance of a number of patterns characterising several real-word hypergraphs.
+Methods
+Formalism and basic quantities
 A hypergraph can be defined as a pair $H(\mathcal{V},\mathcal{E}_{H})$ where $\mathcal{V}$ is the set of vertices and $\mathcal{E}_{H}$ is the set of hyperedges.
-Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in Berge (1967, 1973), where hyperedges ‘link’ one or more vertices, and in Voloshin (2009), where hyperedges are allowed to be empty sets as well.
-Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: in words, we consider non-simple hypergraphs, admitting loops and parallel edges (i.e.
-hyperedges involving exactly the same set nodes) of any size, including 0 (corresponding to empty hyperedges), and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
-$\mathbf{I}\>=\>\begin{tabular}[]{c|c|c|c|c|c|}&$\mathbf{e_{1}}$&$\mathbf{e_{2}}$&$\mathbf{e_{3}}$&$\mathbf{e_{4}}$&$\mathbf{e_{5}}$\\ \hline\cr$\mathbf{n_{1}}$&0&1&1&0&1\\ \hline\cr$\mathbf{n_{2}}$&0&0&1&1&1\\ \hline\cr$\mathbf{n_{3}}$&0&1&1&1&0\\ \hline\cr$\mathbf{n_{4}}$&1&0&0&1&1\\ \hline\cr$\mathbf{n_{5}}$&1&1&1&1&0\\ \hline\cr$\mathbf{n_{6}}$&0&1&0&0&0\\ \hline\cr\end{tabular}$
-Table 1: Example of incidence matrix describing a binary, undirected hypergraph with $6$ nodes and $5$ hyperedges.
-The generic entry $I_{i\alpha}$ is 1 if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise (see also Fig.
-1).
+Moving from the observation that the edge set $\mathcal{E}_{G}$ of a traditional, binary, undirected graph $G(\mathcal{V},\mathcal{E}_{G})$ is a subset of the power set of $\mathcal{V}$, several definitions of the hyperedge set $\mathcal{E}_{H}$ have been provided: the two most popular ones are those proposed in [29, 30], where hyperedges tie one or more vertices, and in [31], where hyperedges are allowed to be empty sets as well.
+Hereby, we adopt the definition according to which $\mathcal{E}_{H}$ is a multiset of the power set of $\mathcal{V}$: since the concept of ‘multiset’ extends the concept of ‘set’, allowing for multiple instances of (each of) its elements, our choice implies that we are considering non-simple hypergraphs, admitting loops and parallel edges (i.e.
+hyperedges involving exactly the same nodes) of any size, including 0 (corresponding to empty hyperedges) and $|\mathcal{V}|$ (corresponding to hyperedges clustering all vertices together).
 As for traditional graphs, an algebraic representation of hypergraphs can be devised as well.
-For a formal analogy with the traditional case, let us call the cardinality of the set of nodes $|\mathcal{V}|=N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|=L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge.
+In analogy with the traditional case, we call the cardinality of the set of nodes $|\mathcal{V}|\equiv N$ and the cardinality of the set of hyperedges $|\mathcal{E}_{H}|\equiv L$: then, we consider the $N\times L$ table known as incidence matrix, each row of which corresponds to a node and each column of which corresponds to a hyperedge.
 If we indicate the incidence matrix with $\mathbf{I}$, its generic entry $I_{i\alpha}$ will be $1$ if vertex $i$ belongs to hyperedge $\alpha$ and 00 otherwise.
 Notice that the number of $1$s along each row can vary between 00 and $L$, the former case indicating an isolated node and the latter one indicating a node that belongs to each hyperedge; similarly, the number of $1$s along each column can vary between 00 and $N$, the former case indicating an empty hyperedge and the latter one indicating a hyperedge that includes all nodes.
-As explicitly noticed elsewhere Nakajima et al.
-(2021); Barthelemy (2022), representing a hypergraph via its incidence matrix is equivalent at considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$.
-Table 1 represents the incidence matrix describing the binary, undirected hypergraph shown in Fig.
-1.
+As explicitly noticed elsewhere [9, 10, 12], representing a hypergraph via its incidence matrix is equivalent to considering the bipartite graph defined by the sets $\mathcal{V}$ and $\mathcal{E}_{H}$ more formally, the function that assigns each hypergraph to the bipartite graph associated with it is a bijection when both nodes and hyperedges are uniquely labeled [9].
+For instance, the incidence matrix $\mathbf{I}$ describing the binary, undirected hypergraph shown in Fig.
+1 is the following:
+Figure 1: Cartoon representation of a hypergraph. Black dots represent nodes, while the colored shapes represent hyperedges. The incidence matrix describing the present hypergraph is defined by Eq. 1.
+[…]
 Once the incidence matrix has been defined, several quantities needed for the description of hypergraphs can be defined quite straightforwardly: for example, the ‘degree of node $i$’ (hereby, degree) reads
@@
 $\sum_{i=1}^{N}k_{i}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}=\sum_{\alpha=1}^{L}\sum_{i=1}^{N}I_{i\alpha}=\sum_{\alpha=1}^{L}h_{\alpha}\equiv T$.
-Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information on the hyperedge size.
-Figure 1: Cartoon representation of the hypergraph whose incidence matrix is shown in Table 1.
-III III. Binary, undirected hypergraphs randomisation
-An early attempt to define randomisation models for hypergraphs can be found in Ghoshal et al.
-(2009).
-Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network Nakajima et al.
-(2021).
-Considering the incidence matrix has two, clear advantages over the tensor-based representation employed in Ghoshal et al.
-(2009); Nakajima et al.
-(2021): i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
-In order to extend the rich set of null models induced by graph-specific, global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting.
+Importantly, a node degree no longer coincides with the number of its neighbours: instead, it matches the number of hyperedges it belongs to; a hyperdegree, instead, provides information about the hyperedge size.
+Analogously, $T$ paves the way for the alternative definition of ‘density of connections’ reading $\rho=T/NL\equiv h/N$, i.e.
+the ratio between the (average) number of nodes each hyperedge clusters and the total number of nodes.
+Binary, undirected hypergraphs randomisation
+An early attempt to define randomisation algorithms for hypergraphs within a framework closely resembling ours can be found in [32].
+Its authors, however, have just considered hyperedges that are incident to triples of nodes - a framework that has been, later, applied to the study of the World Trade Network [10].
+Considering the incidence matrix has two clear advantages over the tensor-based representation employed in [32, 10]: i) generality, because the incidence matrix allows hyperedges of any size to be handled at once; ii) compactness, because the order of the tensor $\mathbf{I}$ never exceeds two, hence allowing any hypergraph to be represented as a traditional, bipartite graph.
+In order to extend the rich set of null models induced by graph-specific global and local constraints to hypergraphs, we, first, need to identify the quantities that can play this role within the novel setting.
 In what follows, we will consider the total number of $1$s, i.e.
@@
 $\{k_{i}\}_{i=1}^{N}$ and $\{h_{\alpha}\}_{\alpha=1}^{L}$ either separately or in a joint fashion; moreover, we will distinguish between microcanonical and canonical randomisation techniques.
-III.1 A. Homogeneous benchmarks:
+Homogeneous benchmarks:
 the Random Hypergraph Model (RHM)
-The model is defined by just one, global constraint, that, in our case, reads $T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}$.
-Its microcanonical version extends the model by Erdös and Rényi Erdös and Rényi (1959) - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of 1s, say $T^{*}$: they are
+Microcanonical formulation
+The model is defined by just one, global constraint, that, in our case, reads
+$T=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}I_{i\alpha};$ (4)
+its microcanonical version extends the model by Erdös and Rényi [33] - also known as Random Graph Model (RGM) - to hypergraphs and prescribes to count the number of incidence matrices that are compatible with a given, total number of $1$s, say $T^{*}$: they are
 $\Omega_{\text{RHM}}=\binom{V}{T^{*}}$ (5)
-with $V=NL$ being the total number of entries of the incidence matrix $\mathbf{I}$.
-Once the total number of configurations composing the microcanonical ensemble is determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of 1s is kept fixed while any other correlation is destroyed.
-The canonical version of the RHM, instead, extends the model by Gilbert Gilbert (1959) and rests upon the constrained maximisation of Shannon entropy, i.e.
+with $V\equiv NL$ being the total number of entries of the incidence matrix $\mathbf{I}$.
+Once the total number of configurations composing the microcanonical ensemble has been determined, a procedure to generate them is needed: in the case of the RHM, it simply boils down to reshuffling the entries of the incidence matrix, a procedure ensuring that the total number of $1$s is kept fixed while any, other correlation is destroyed.
+Canonical formulation
+The canonical version of the RHM, instead, extends the model by Gilbert [34] and rests upon the constrained maximisation of Shannon entropy, i.e.
 $\mathscr{L}\equiv S[P]-\sum_{i=0}^{M}\theta_{i}[P(\mathbf{I})C_{i}(\mathbf{I})-\langle C_{i}\rangle]$ (6)
-where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties.
+where $S[P]=-\sum_{\mathbf{I}\in\mathcal{I}}P(\mathbf{I})\ln P(\mathbf{I})$, $C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper topological properties.
 The sum defining Shannon entropy runs over the set $\mathcal{I}$ of incidence matrices described in the introductory paragraph and known as canonical ensemble.
@@
 $P(\mathbf{I})=\frac{e^{-H(\mathbf{I})}}{Z}=\frac{e^{-H(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-H(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{M}\theta_{i}C_{i}(\mathbf{I})}}.$ (7)
-In the simplest case, the only, global constraint is represented by $T$ and leads to the expression
-$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}$
-$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}$
-$=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (6)
+In the simplest case, the only global constraint is represented by $T$ and leads to the expression
+$P(\mathbf{I})$ $=\frac{e^{-\theta T(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\theta T(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\theta I_{i\alpha}}}=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}x^{I_{i\alpha}}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x)^{-1}$ (8)
 that can be re-written as
@@
 with $e^{-\theta}\equiv x$ and $p\equiv x/(1+x)$.
-The canonical ensemble, now, includes all $N\times L$ rectangular matrices whose number of entries equating 1 ranges from 0 to $NL$.
+The canonical ensemble, now, includes all $N\times L$, rectangular matrices whose number of entries equal to 1 ranges from 0 to $NL$.
 According to such a model, the entries of the incidence matrix are i.i.d.
 Bernoulli random variables, i.e.
-$I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$: as a consequence, the total number of 1s, the degrees and the hyperdegrees obey Binomial distributions - being all defined as sums of i.i.d.
-Bernoulli random variables.
-Specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
+$I_{i\alpha}\sim\text{Ber}(p)$, $\forall\>i,\alpha$; as a consequence, the total number of $1$s, the degrees and the hyperdegrees obey Binomial distributions, being all defined as sums of i.i.d.
+Bernoulli random variables: specifically, $T\sim\text{Bin}(NL,p)$, $k_{i}\sim\text{Bin}(L,p)$, $\forall\>i$ and $h_{\alpha}\sim\text{Bin}(N,p)$, $\forall\>\alpha$, in turn, implying that $\langle T\rangle_{\text{RHM}}=NLp$, $\langle k_{i}\rangle_{\text{RHM}}=Lp$, $\forall\>i$ and $\langle h_{\alpha}\rangle_{\text{RHM}}=Np$, $\forall\>\alpha$.
+Parameter estimation
 In order to ensure that $\langle T\rangle_{\text{RHM}}=T^{*}$, parameters have to be tuned opportunely.
-To this aim, the likelihood maximisation principle can be invoked Garlaschelli and Loffredo (2008): it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter(s) that define it.
-Such a recipe leads us to find $p=T^{*}/NL$, with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
-The RHM (also considered in Barthelemy (2022), although a derivation from first principles is missing, there) is formally equivalent to the Bipartite Random Graph Model (BiRGM) Saracco et al.
-(2015).
+To this aim, the likelihood maximisation principle can be invoked [19]: it prescribes to maximise the function $\mathcal{L}(\theta)\equiv\ln P(\mathbf{I}^{*}|\theta)$ with respect to the unknown parameter that defines it.
+Such a recipe leads us to find
+$p=\rho^{*}=\frac{T^{*}}{NL}$ (10)
+with $T^{*}=T(\mathbf{I}^{*})$ indicating the empirical value of the constraint defining the RHM.
+The RHM (also considered in [12], although without providing any derivation from first principles, and in [16], although without providing any recipe for the estimation of its parameter) is formally equivalent to the Bipartite Random Graph Model (BiRGM) [35].
 Such an identification is guaranteed by our focus on non-simple hypergraphs.
-III.1.1 1. Estimation of the number of empty hyperedges
-As non-simple hypergraphs admit the presence of empty as well as parallel hyperedges, let us evaluate how frequently they appear in the ensembles induced by our benchmarks.
-To this aim, let us start by considering the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero.
-Upon remembering that $h_{\alpha}\sim\text{Bin}(N,p)$, one finds
-$p_{0}\equiv(1-p)^{N}.$ (8)
-Such a quantity leads us to find the probability of observing at least one, empty hyperedge (or, in other terms, the complementary of the probability that no hyperedge is empty), i.e.
-$p_{1}\equiv 1-(1-p_{0})^{L}=1-[1-(1-p)^{N}]^{L},$ (9)
-and the expected number of empty hyperedges: upon remembering that they are i.i.d. Binomial random variables, one finds
-$\langle N_{0}\rangle\equiv Lp_{0}=L(1-p)^{N}.$ (10)
-Let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the RHM, as the density of 1s in the incidence matrix, i.e.
+Estimation of the number of empty hyperedges
+Non-simple hypergraphs admit the presence of empty as well as parallel hyperedges.
+As this type of structures is associated with configurations that may be regarded as problematic (since not observed in empirical data), we evaluate how frequently they appear in the ensembles induced by our benchmarks.
+Let us denote the number of empty hyperedges, i.e.
+the number of hyperedges whose hyperdegree equals zero, with $N_{\emptyset}$: since the hyperdegrees are i.i.d.
+Binomial random variables, $N_{\emptyset}\sim\text{Bin}(L,p_{\emptyset})$ where
+${\color[rgb]{0,0,0}p_{\emptyset}\equiv(1-p)^{N}}$ (11)
+is the probability for the generic hyperedge to be empty or, equivalently, for its hyperdegree to equal zero; the expected number of empty hyperedges reads
+${\color[rgb]{0,0,0}\langle N_{\emptyset}\rangle=Lp_{\emptyset}=L(1-p)^{N}.}$ (12)
+Let us, now, inspect the behaviour of $p_{\emptyset}$ on the ensemble induced by the RHM as the density of $1$s in the incidence matrix, i.e.
 $p=T/NL$, varies.
@@
 In the dense case, one finds
-$\lim_{T\to NL}p_{0}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (11)
-a relationship inducing $p_{1}\xrightarrow[]{T\to NL}0$ and $\langle N_{0}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases.
+$\lim_{T\to NL}p_{\emptyset}$ $=\lim_{T\to NL}\left(1-\frac{T}{NL}\right)^{N}=0,$ (13)
+a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{T\to NL}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of $1$s increases.
 Consistently, in the sparse case one finds
-$\lim_{T\to 0}p_{0}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (12)
-a relationship inducing $p_{1}\xrightarrow[]{T\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
-To evaluate the density of 1s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e.
-$p=O(1/N)$.
-Since $p=T/NL$ and $h\equiv T/L$ remains finite, the probability for a generic hyperedge to be empty remains finite as well: consistently, the expected number of empty hyperedges becomes
-$\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}L\left(1-\frac{h}{N}\right)^{N}=Le^{-h};$ (13)
+$\lim_{T\to 0}p_{\emptyset}=\lim_{T\to 0}\left(1-\frac{T}{NL}\right)^{N}=1,$ (14)
+a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{T\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of $1$s decreases.
+To evaluate the density of $1$s in the incidence matrix in correspondence of which the transition from the sparse to the dense regime happens, let us consider the case $N\gg 1$: more formally, this amounts to consider the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$ equivalently, defined by posing $p=O(1/N)$.
+Since $p=T/NL$ and $h=T/L$ remains finite, the probability for the generic hyperedge to be empty obeys the relationship
+${\color[rgb]{0,0,0}\lim_{N\to+\infty}p_{\emptyset}=\lim_{N\to+\infty}\left(1-\frac{h}{N}\right)^{N}=e^{-h},}$ (15)
+i.e. remains finite as well: consistently, the expected number of empty hyperedges becomes
+${\color[rgb]{0,0,0}\lim_{N\to+\infty}\langle N_{\emptyset}\rangle=\lim_{N\to+\infty}Lp_{\emptyset}=Le^{-h};}$ (16)
 upon imposing $Le^{-h}\leq 1$, i.e.
-that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}=\ln L$.
-In words, a value $p>p_{f}^{\text{RHM}}\equiv h_{f}^{\text{RHM}}/N=\ln L/N$ ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one.
-The calculation above complements the result derived in Barthelemy (2022) about the percolation threshold for hypergraphs (see also Appendix A).
-III.1.2 2. Estimation of the number of parallel hyperedges
+that the expected number of empty hyperedges is at most 1, one derives what may be called filling threshold, corresponding to $h_{f}^{\text{RHM}}\equiv\ln L$.
+In words, a value
+${\color[rgb]{0,0,0}p>p_{f}^{\text{RHM}}=\frac{h_{f}^{\text{RHM}}}{N}=\frac{\ln L}{N}}$ (17)
+ensures that the expected number of empty hyperedges in our random hypergraph is strictly less than one.
+As a last observation, let us notice that evaluating $p_{\emptyset}$ in correspondence of the filling threshold returns the value $1/L$.
+A comparison with simple graphs: estimating the number of isolated nodes
+A similar line of reasoning can be repeated for traditional graphs, the aim being, now, that of estimating $N_{0}$, i.e.
+the number of isolated nodes.
+To this aim, let us consider the asymptotic framework defined by letting $N\to+\infty$ while posing $L=O(N)$ equivalently, defined by posing $q=O(1/N)$.
+Since $q=2L/N(N-1)$ and $k\equiv 2L/(N-1)$ remains finite, the probability for the generic node $i$ to be isolated obeys the relationship
+${\color[rgb]{0,0,0}\lim_{N\to+\infty}q_{0}=\lim_{N\to+\infty}(1-q)^{N-1}=\lim_{N\to+\infty}\left(1-\frac{k}{N-1}\right)^{N-1}=e^{-k},}$ (18)
+i.e. remains finite as well: this, in turn, implies that the expected number of isolated nodes $\langle N_{0}\rangle=Nq_{0}$ obeys the relationship
+${\color[rgb]{0,0,0}\lim_{N\to+\infty}\langle N_{0}\rangle=\lim_{N\to+\infty}Nq_{0}=Ne^{-k};}$ (19)
+upon imposing $Ne^{-k}\leq 1$, i.e. that the expected number of isolated nodes is at most 1, one derives the connectivity threshold, corresponding to $k_{c}^{\text{RGM}}\equiv\ln N$. In words, a value
+${\color[rgb]{0,0,0}q>q_{c}^{\text{RGM}}=\frac{k_{c}^{\text{RGM}}}{N}=\frac{\ln N}{N}}$ (20)
+ensures that the expected number of isolated nodes in our random graph is strictly less than one [1].
+As a last observation, let us notice that evaluating $q_{0}$ in correspondence of the connectivity threshold returns the value $1/N$.
+We also note that $N$ is the only quantity playing a relevant role in the case of graphs, while an interplay between $L$ and $N$ can be observed in the case of hypergraphs: in both cases, however, a condition on connectivity is present, driven by the request that the objects under investigations (nodes on the one hand and hyperedges on the other) have a non-zero number of connections.
+Estimation of the number of parallel hyperedges
 Let us, now, move to considering the issue of parallel hyperedges.
@@
 $d_{\alpha\beta}\equiv\sum_{i=1}^{N}[I_{i\alpha}(1-I_{i\beta})+I_{i\beta}(1-I_{i\alpha})],$ (21)
-a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$.
-Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that $P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ and $\langle d_{\alpha\beta}\rangle=2p(1-p)N$, $\forall\>\alpha\neq\beta$.
-Notice that
-$\lim_{T\to NL}P(d_{\alpha\beta}=0)$ $=\lim_{T\to 0}P(d_{\alpha\beta}=0)=1$ (15)
-and
-$\lim_{T\to NL}\langle d_{\alpha\beta}\rangle$ $=\lim_{T\to 0}\langle d_{\alpha\beta}\rangle=0$ (16)
-as $(2T/NL)(1-T/NL)\to 0$ in both regimes: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s.
+a sum whose generic addendum is 1 in just two cases: either $I_{i\alpha}=1$ and $I_{i\beta}=0$ or $I_{i\alpha}=0$ and $I_{i\beta}=1$. Since $d_{\alpha\beta}\sim\text{Bin}(N,2p(1-p))$, one finds that
+$p_{\mathbin{/\!/}}^{\alpha\beta}\equiv P(d_{\alpha\beta}=0)=[1-2p(1-p)]^{N}$ (22)
+and that
+$\langle d_{\alpha\beta}\rangle=2p(1-p)N$ (23)
+$\forall\>\alpha\neq\beta$. Since $p(1-p)=(T/NL)(1-T/NL)$, one finds that
+$\lim_{T\to NL}p(1-p)$ $=\lim_{T\to 0}p(1-p)=0,$ (24)
+i.e.
+$p(1-p)$ vanishes in both regimes, a result further implying that both $P(d_{\alpha\beta}=0)\xrightarrow[]{T\to NL}1$ and $P(d_{\alpha\beta}=0)\xrightarrow[]{T\to 0}1$ and that both $\langle d_{\alpha\beta}\rangle\xrightarrow[]{T\to NL}0$ and $\langle d_{\alpha\beta}\rangle\xrightarrow[]{T\to 0}0$: in words, the probability of observing parallel hyperedges progressively rises both as a consequence of having many $1$s and as a consequence of having few $1$s.
 Analogously for the expected Hamming distance.
-Let us, now, evaluate the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$, within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$, i.e.
-$p=O(1/N)$.
+Let us, now, evaluate the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ within the asymptotic framework defined by letting $N\to+\infty$ while posing $T=O(L)$ equivalently, defined by posing $p=O(1/N)$.
 Since $p=T/NL$ and $h\equiv T/L$ remains finite, one finds that
@@
 upon imposing $2h\geq 1$, i.e.
-that the expected Hamming distance between any, two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}=1/2$.
-In words, a value $p>p_{r}^{\text{RHM}}\equiv h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element (see also Appendix A).
-III.1.3 3. Estimation of the overlap between hyperedges
-Finally, let us move to considering the issue of overlapping hyperedges. To this aim, let us require that the probability for any, two hyperedges to overlap amounts to 1. In formulas,
-$1-(1-p^{2})^{N}=1,$ (18)
-a relationship leading to the equation
-$(1-p^{2})^{N}\simeq 1-Np^{2}=0$ (19)
-that holds true in the sparse regime and further implies
-$h^{2}=N;$ (20)
-such a value, firstly derived in Barthelemy (2022), individuates the so-called percolation threshold, corresponding to $h_{p}^{\text{RHM}}=\sqrt{N}$.
-In words, a value $p>p_{p}^{\text{RHM}}\equiv h_{p}^{\text{RHM}}/N=1/\sqrt{N}$ ensures that any, two hyperedges in our random hypergraph share at least one node.
-Since $h_{r}^{\text{RHM}}\leq h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, such a result contributes to describe the following situation: by progressively rising the parameter $p$, three, different thresholds are met, i.e.
-i) the resolution threshold $p_{r}^{\text{RHM}}=1/2N$, ensuring that, on average, any, two hyperedges differ by at least one element; ii) the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, ensuring that the expected number of empty hyperedges is strictly less than one; iii) the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{N}$, ensuring that any, two hyperedges overlap.
-III.2 B. Heterogeneous benchmarks:
+that the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ is at least 1, one derives what may be called resolution threshold, corresponding to $h_{r}^{\text{RHM}}\equiv 1/2$.
+In words, a value $p>p_{r}^{\text{RHM}}=h_{r}^{\text{RHM}}/N=1/2N$ ensures that, on average, any, two hyperedges $\alpha$ and $\beta$ differ by at least one element.
+To derive a global condition on the total number of parallel hyperedges, note that, although the overlaps between pairs of hyperedges cannot be treated as i.i.d.
+random variables, the expected number of parallel hyperedges can still be computed explicitly.
+Upon posing $p_{\mathbin{/\!/}}\equiv p_{\mathbin{/\!/}}^{\alpha\beta}$, it reads
+${\color[rgb]{0,0,0}\langle N_{\mathbin{/\!/}}\rangle=\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}=\frac{L(L-1)}{2}p_{\mathbin{/\!/}}.}$ (26)
+Considering that $p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$ and imposing $\langle N_{\mathbin{/\!/}}\rangle\leq 1$, i.e.
+that the expected number of parallel hyperedges is at most 1, one derives what may be called a multiple resolution threshold, corresponding to $h_{m}^{\text{RHM}}\equiv\ln L-\ln\sqrt{2}\lesssim h_{f}^{\text{RHM}}$.
+In words, a value
+${\color[rgb]{0,0,0}p>p_{m}^{\text{RHM}}=\frac{h_{m}^{\text{RHM}}}{N}=\frac{\ln L}{N}-\frac{\ln\sqrt{2}}{N}\lesssim\frac{\ln L}{N}=\frac{h_{f}^{\text{RHM}}}{N}=p_{f}^{\text{RHM}}}$ (27)
+(also) ensures that the expected number of parallel hyperedges in our random hypergraph is strictly less than one.
+Estimation of the percolation threshold
+The two thresholds derived in the previous subsections emerge in consequence of the attempts of solving the problems related to the appearance of empty as well as parallel hyperedges.
+Remarkably, a third threshold exists: known as percolation threshold, it was, first, derived in [12], following the definition according to which any two hyperedges are said to be connected if they share at least one node.
+Here, we re-derive the percolation threshold by considering the ‘hypergraph to graph’ projection (see also below): more formally, the total number of nodes shared by hyperedge $\alpha$ with any other hyperedge reads
+${\color[rgb]{0,0,0}\sigma_{\alpha}=\sum_{i=1}^{N}\sum_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}I_{i\alpha}I_{i\beta},}$ (28)
+its expected value being
+${\color[rgb]{0,0,0}\langle\sigma_{\alpha}\rangle=N(L-1)p^{2}\simeq NLp^{2};}$ (29)
+imposing $\langle\sigma_{\alpha}\rangle=1$ leads to find the value
+${\color[rgb]{0,0,0}p_{p}^{\text{RHM}}=\frac{h_{p}^{\text{RHM}}}{N}=\frac{1}{\sqrt{NL}}}$ (30)
+which, in turn, induces the value $h_{p}^{\text{RHM}}\equiv\sqrt{N/L}$.
+In words, a value $p>p_{p}^{\text{RHM}}$ ensures that any two hyperedges in our random hypergraph share, on average, at least one node.
+Heterogeneous benchmarks:
 the Hypergraph Configuration Model (HCM)
+Microcanonical formulation
 The number of constraints can be enlarged to include the degrees, i.e.
@@
 the sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$.
-Counting the number of configurations on which both sequences match their empirical values is a hard task - although numerical recipes to shuffle the entries of a rectangular matrix, while preserving its marginals, exist Strona et al.
-(2014); Carstens (2015); Chodrow (2020).
-Solving the corresponding problem in the canonical framework is, instead, straightforward.
-Indeed, Shannon entropy maximisation leads to
-$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}$
-$=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
+Although counting the number of configurations on which both sequences match their empirical values is a hard task, numerical recipes that shuffle the entries of a rectangular matrix, while preserving its marginals, exist [36, 37, 9, 38].
+It should be, however, noticed that, if not carefully implemented, algorithms of the kind may lead to a non-uniform exploration of the space of configurations [39, 40]; moreover, the issue concerning the time needed to collect a sufficiently large number of configurations should be addressed even in presence of an ergodic system.
+A recent proposal is that of extending the traditional Curveball algorithm to hypergraphs [38].
+Canonical formulation
+Solving the corresponding problem in the canonical framework is, instead, straightforward. Indeed, Shannon entropy maximisation leads to
+$P(\mathbf{I})$ $=\frac{e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\alpha_{i}k_{i}(\mathbf{I})-\sum_{\alpha=1}^{L}\beta_{\alpha}h_{\alpha}(\mathbf{I})}}=\frac{e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}{\sum_{\mathbf{I}\in\mathcal{I}}e^{-\sum_{i=1}^{N}\sum_{\alpha=1}^{L}(\alpha_{i}+\beta_{\alpha})I_{i\alpha}}}$
 $=\prod_{i=1}^{N}x_{i}^{k_{i}(\mathbf{I})}\prod_{\alpha=1}^{L}y_{\alpha}^{h_{\alpha}(\mathbf{I})}\prod_{i=1}^{N}\prod_{\alpha=1}^{L}(1+x_{i}y_{\alpha})^{-1},$ (31)
@@
 As a consequence, both degrees and hyperdegrees obey Poisson-Binomial distributions, i.e.
-$k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ Saracco et al.
-(2017).
+$k_{i}\sim\text{PoissBin}(L,\{p_{i\alpha}\}_{\alpha=1}^{L})$, $\forall\>i$ and $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, $\forall\>\alpha$ [24].
+Parameter estimation
 In this case, solving the likelihood maximisation problem amounts to solve the system of coupled equations
@@
 $p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (35)
-The HCM reduces to a ‘partial’ Configuration Model Saracco et al.
-(2017) when either the degree or the hyperdegree sequence is left unconstrained (see also Appendix B).
-The canonical ensemble of each randomisation model (Table 2 in Appendix B sums up the set of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
-The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) Saracco et al. (2015). Such an identification is guaranteed by our focus on non-simple hypergraphs.
-III.2.1 1. Estimation of the number of empty hyperedges
+The HCM reduces to a ‘partial’ Configuration Model [24] when either the degree or the hyperdegree sequence is left unconstrained (see also Section A in the Appendix).
+The canonical ensemble of each randomisation model (Table 1 in the Appendix sums up the sets of constraints defining them) can be explicitly sampled by considering each entry of $\mathbf{I}$, drawing a real number $u_{i\alpha}\in U[0,1]$ and posing $I_{i\alpha}=1$ if $u_{i\alpha}\leq p_{i\alpha}$, $\forall\>i,\alpha$.
+The HCM is formally equivalent to the Bipartite Configuration Model (BiCM) [35]. Such an identification is guaranteed by our focus on non-simple hypergraphs.
+Estimation of the number of empty hyperedges
 Let us, now, consider the probability for the generic hyperedge $\alpha$ to be empty or, in other terms, that its hyperdegree $h_{\alpha}$ is zero.
 Upon remembering that $h_{\alpha}\sim\text{PoissBin}(N,\{p_{i\alpha}\}_{i=1}^{N})$, one finds
-$p_{0}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha}).$ (26)
-Such a quantity leads to the probability of observing at least one, empty hyperedge, now reading
-$p_{1}\equiv 1-\prod_{\alpha=1}^{L}\left[1-p_{0}^{\alpha}\right]=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right],$ (27)
-and to the expected number of empty hyperedges
-$\langle N_{0}\rangle\equiv\sum_{\alpha=1}^{L}p_{0}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (28)
-As in the previous section, let us, now, inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM, as the density of 1s in the incidence matrix varies.
-Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can still be drawn within a simpler framework.
+$p_{\emptyset}^{\alpha}\equiv\prod_{i=1}^{N}(1-p_{i\alpha})$ (36)
+while the expected number of empty hyperedges, now, reads
+$\langle N_{\emptyset}\rangle\equiv\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}=\sum_{\alpha=1}^{L}\prod_{i=1}^{N}(1-p_{i\alpha}).$ (37)
+As previously done, let us inspect the behaviour of the aforementioned quantities on the ensemble induced by the HCM as the density of $1$s in the incidence matrix varies.
+Although it depends on (the heterogeneity of) the sets of coefficients $\{x_{i}\}_{i=1}^{N}$ and $\{y_{\alpha}\}_{\alpha=1}^{L}$, general conclusions can be still drawn within a simpler framework.
 To this aim, let us consider the functional form reading
 $p_{i\alpha}=\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}},\>\forall\>i,\alpha$ (38)
-where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of 1s in the incidence matrix [^1] Partial Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$..
-Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution.
+where the vector of fitnesses $\{f_{i}\}_{i=1}^{N}$ accounts for the heterogeneity of nodes, the vector of fitnesses $\{g_{\alpha}\}_{\alpha=1}^{L}$ accounts for the heterogeneity of hyperedges and $z$ tunes the density of $1$s in the incidence matrix - ‘partial’ Configuration Models are recovered upon posing either $f_{i}=1$, $\forall\>i$ or $g_{\alpha}=1$, $\forall\>\alpha$.) Within such a framework, the fitnesses of the nodes and the fitnesses of the hyperedges can be drawn from any distribution.
 The dense and sparse regimes are, now, defined by the positions $z\to+\infty$ and $z\to 0$, respectively.
 In the dense case, one finds
-$\lim_{z\to+\infty}p_{0}^{\alpha}$ $=\lim_{z\to+\infty}\prod_{i=1}^{N}(1-p_{i\alpha})$
-$=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (30)
-a relationship inducing $p_{1}\xrightarrow[]{z\to+\infty}0$ and $\langle N_{0}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of 1s increases.
+$\lim_{z\to+\infty}p_{\emptyset}^{\alpha}=\lim_{z\to+\infty}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=0,$ (39)
+a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{z\to+\infty}0$: in words, the probability of observing empty hyperedges progressively vanishes as the density of $1$s increases.
 Consistently, in the sparse case one finds
-$\lim_{z\to 0}p_{0}^{\alpha}$ $=\lim_{z\to 0}\prod_{i=1}^{N}(1-p_{i\alpha})$
-$=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (31)
-a relationship inducing $p_{1}\xrightarrow[]{z\to 0}1$ and $\langle N_{0}\rangle\xrightarrow[]{z\to 0}L$.
-In words, the probability of observing empty hyperedges progressively rises as the density of 1s decreases.
-III.2.2 2. Estimation of the number of parallel hyperedges
-Let us, now, focus on the issue of parallel hyperedges.
+$\lim_{z\to 0}p_{\emptyset}^{\alpha}=\lim_{z\to 0}\prod_{i=1}^{N}\left(1-\frac{zf_{i}g_{\alpha}}{1+zf_{i}g_{\alpha}}\right)=1,$ (40)
+a relationship inducing $\langle N_{\emptyset}\rangle\xrightarrow[]{z\to 0}L$: in words, the probability of observing empty hyperedges progressively rises as the density of $1$s decreases.
+For what concerns the filling threshold, a derivation that is similar-in-spirit to the one carried out for the case of the RHM can be sketched.
+Let us pose ourselves in the sparse regime: since $1-p_{i\alpha}\simeq e^{-p_{i\alpha}}$, the probability for the generic hyperedge to be empty satisfies the chain of relationships
+${\color[rgb]{0,0,0}p_{\emptyset}^{\alpha}=\prod_{i=1}^{N}(1-p_{i\alpha})\simeq\prod_{i=1}^{N}e^{-p_{i\alpha}}=e^{-\sum_{i=1}^{N}p_{i\alpha}}=e^{-h_{\alpha}};}$ (41)
+consistently, the expected number of empty hyperedges becomes
+${\color[rgb]{0,0,0}\langle N_{\emptyset}\rangle=\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}\simeq\sum_{\alpha=1}^{L}e^{-h_{\alpha}}}$ (42)
+and imposing $\langle N_{\emptyset}\rangle\leq 1$, i.e.
+that the expected number of empty hyperedges is at most 1, one derives a global condition to be satisfied by the hyperdegrees.
+In general terms, the aforementioned condition leads to require $e^{-h_{\alpha}}=O(1/L)$, i.e.
+$h_{\alpha}=O(\ln L)$, $\forall\>\alpha$ and $p_{i\alpha}=O(\ln L/N)$, $\forall\>i,\alpha$.
+A comparison with simple graphs: estimating the number of isolated nodes
+Coming to traditional graphs, the aim is, now, to estimate the number of isolated nodes. In this case, $1-p_{ij}\simeq e^{-p_{ij}}$ and the probability for the generic node $i$ to be isolated reads
+${\color[rgb]{0,0,0}q_{0}^{i}=\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}(1-p_{ij})\simeq\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}e^{-p_{ij}}=e^{-\sum_{j(\neq i)=1}^{N}p_{ij}}=e^{-k_{i}};}$ (43)
+consistently, the expected number of isolated nodes becomes
+${\color[rgb]{0,0,0}\langle N_{0}\rangle=\sum_{i=1}^{N}q_{0}^{i}\simeq\sum_{i=1}^{N}e^{-k_{i}}}$ (44)
+and imposing $\langle N_{0}\rangle\leq 1$, i.e.
+that the expected number of isolated nodes is at most 1, one derives a global condition to be satisfied by the degrees.
+In general terms, the aforementioned condition leads to require $e^{-k_{i}}=O(1/N)$, i.e.
+$k_{i}=O(\ln N)$, $\forall\>i$ and $p_{ij}=O(\ln N/N)$, $\forall\>i<j$.
+Estimation of the number of parallel hyperedges
 As in the case of the RHM, we consider the Hamming distance between the columns representing the two hyperedges $\alpha$ and $\beta$.
-Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ and $\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$, $\forall\>\alpha\neq\beta$.
-Notice that
-$\lim_{z\to+\infty}P(d_{\alpha\beta}=0)$ $=\lim_{z\to 0}P(d_{\alpha\beta}=0)=1$ (32)
+Since, now, $d_{\alpha\beta}\sim\text{PoissBin}(N,\{q_{i}^{\alpha\beta}\}_{i=1}^{N})$, where $q_{i}^{\alpha\beta}\equiv p_{i\alpha}(1-p_{i\beta})+p_{i\beta}(1-p_{i\alpha})$ with $p_{i\alpha}=zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ and $p_{i\beta}=zf_{i}g_{\beta}/(1+zf_{i}g_{\beta})$, one finds that
+$P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$ (45)
+and that
+$\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}$ (46)
+$\forall\>\alpha\neq\beta$. Since
+$\lim_{z\to+\infty}q_{i}^{\alpha\beta}$ $=\lim_{z\to 0}q_{i}^{\alpha\beta}=0,$ (47)
+i.e.
+$q_{i}^{\alpha\beta}$ vanishes in both regimes, one finds that both $P(d_{\alpha\beta}=0)\xrightarrow[]{z\to+\infty}1$ and $P(d_{\alpha\beta}=0)\xrightarrow[]{z\to 0}1$ and that both $\langle d_{\alpha\beta}\rangle\xrightarrow[]{z\to+\infty}0$ and $\langle d_{\alpha\beta}\rangle\xrightarrow[]{z\to 0}0$: as in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many $1$s and as a consequence of having few $1$s.
+Analogously for the expected Hamming distance.
+For what concerns the resolution threshold, a derivation that is similar-in-spirit to the one carried out for the case of the RHM can be sketched.
+Let us pose ourselves in the sparse regime and consider that $1-q_{i}^{\alpha\beta}\simeq e^{-q_{i}^{\alpha\beta}}\simeq e^{-(p_{i\alpha}+p_{i\beta})}$.
+As a consequence
+${\color[rgb]{0,0,0}p_{\mathbin{/\!/}}^{\alpha\beta}\equiv P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})\simeq\prod_{i=1}^{N}e^{-q_{i}^{\alpha\beta}}\simeq e^{-\sum_{i=1}^{N}(p_{i\alpha}+p_{i\beta})}=e^{-(h_{\alpha}+h_{\beta})}}$ (48)
 and
-$\lim_{z\to 0}\langle d_{\alpha\beta}\rangle$ $=\lim_{z\to 0}\langle d_{\alpha\beta}\rangle=0$ (33)
-as $q_{i}^{\alpha\beta}\to 0$ in both regimes.
-As in the case of the RHM, the probability of observing parallel hyperedges progressively rises both as a consequence of having many 1s and as a consequence of having few 1s.
-Analogously for the expected Hamming distance.
-When considering heterogeneous benchmarks, closed-form expressions for the asymptotic estimates can be provided only in specific cases (see also Appendix B).
-Figure 2: Left panel: trend of $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e.
-the probability of observing at least one empty hyperedge as a function of $T/NL$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (whose related, critical density value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ individuates a steep transition separating the sparse regime from the dense one) returns the value $0.63231-(1-1/L)^{L}=$ 0.6323.
-Right panel: trend of $p_{1}$ as defined by Eq.
-(27).
-Notice that the trend induced by the HCM is shifted on the right with respect to the trend induced by the RHM: this is probably due to the presence of small fitnesses increasing the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish.
-The dense (sparse) regime is recovered for large (small) values of $T/NL$.
-Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
-IV IV. Results
-IV.1 A. Hypergraphs in the dense and sparse regime
-Let us start by verifying the correctness of the estimations of the number of empty and parallel hyperedges provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$ (for other values, see Appendix A and Appendix B).
-IV.1.1 1. The Random Hypergraph Model
-Let us start focusing on our homogeneous benchmark.
-The left panel of Fig.
-2 depicts the trend of $p_{1}=1-\left[1-\left(1-\frac{T}{NL}\right)^{N}\right]^{L}$.
-The left panel of Fig.
-3 depicts the trend of $\frac{\langle N_{0}\rangle}{L}=\left(1-\frac{T}{NL}\right)^{N}=p_{0}$ and $P(d_{\alpha\beta}=0)=\left[1-\frac{2T}{NL}\left(1-\frac{T}{NL}\right)\right]^{N}$.
-Each quantity has been plotted as a function of $T/NL\in[10^{-6},1]$.
-The dense (sparse) regime is recovered for large (small) values of $T/NL$.
+${\color[rgb]{0,0,0}\langle d_{\alpha\beta}\rangle=\sum_{i=1}^{N}q_{i}^{\alpha\beta}\simeq\sum_{i=1}^{N}(p_{i\alpha}+p_{i\beta})=h_{\alpha}+h_{\beta}.}$ (49)
+Imposing $\langle d_{\alpha\beta}\rangle\geq 1$, i.e.
+that the expected Hamming distance between any two hyperedges $\alpha$ and $\beta$ is at least 1, amounts to require that $P(d_{\alpha\beta}=0)\leq e^{-1}$ thus recovering the same condition holding true in the case of the RHM where, in fact, $p_{\mathbin{/\!/}}\equiv p_{\mathbin{/\!/}}^{\alpha\beta}\xrightarrow[]{N\to+\infty}e^{-2h}$.
+The expected number of parallel hyperedges, now, reads
+${\color[rgb]{0,0,0}\langle N_{\mathbin{/\!/}}\rangle=\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}^{\alpha\beta}\simeq\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}e^{-(h_{\alpha}+h_{\beta})};}$ (50)
+upon imposing $\langle N_{\mathbin{/\!/}}\rangle\leq 1$, i.e.
+that the expected number of parallel hyperedges is at most 1, one derives a global condition to be satisfied by the hyperdegrees.
+In general terms, the aforementioned condition leads to require $e^{-(h_{\alpha}+h_{\beta})}=O(1/L^{2})$, i.e.
+$h_{\alpha}=O(\ln L)$, $\forall\>\alpha$ and $p_{i\alpha}=O(\ln L/N)$, $\forall\>i,\alpha$.
+Estimation of the percolation threshold
+For what concerns the percolation threshold, the expected value of the total number of nodes shared by hyperedge $\alpha$ with any other hyperedge, now, reads
+${\color[rgb]{0,0,0}\langle\sigma_{\alpha}\rangle=\sum_{i=1}^{N}\sum_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}p_{i\alpha}p_{i\beta}}$ (51)
+and imposing $\langle\sigma_{\alpha}\rangle=1$ leads to a global condition to be satisfied.
+In general terms, the aforementioned condition leads to require $p_{i\alpha}p_{i\beta}=O(1/NL)$, i.e.
+$p_{i\alpha}=O(1/\sqrt{NL})$, $\forall\>i,\alpha$.
+Results
+Hypergraphs in the dense and sparse regime
+Let us start by verifying the correctness of the estimations of the filling, multiple resolution and percolation thresholds provided by our benchmarks: to this aim, we have considered the values $N=300$ and $L=1000$.
+The Random Hypergraph Model
+Each quantity has been plotted as a function of $p\in[10^{-6},1]$.
+The dense (sparse) regime is recovered for large (small) values of $p$.
 Each dot of Figs.
-2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval.
-A first, general comment concerns the agreement between the analytical trends (solid lines) and the numerical estimations (dots), confirming the correctness of our formulas.
-The filling threshold.
-Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{0}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is $p_{1}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$; evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1-(1-1/L)^{L}\simeq 0.6323$ interestingly enough, letting $L$ grow asymptotically as well, the previous value tends to $1-e^{-1}\simeq 0.6321$.
-Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $0.0011/L=$ 0.001: in words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e.
-2/3$\simeq$ 2/3)≃ 2 / 3 ).
-The steepness of the trend of $p_{1}$, however, suggests it to quickly vanish as the density of 1s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
-Figure 3: Left panel: trends of $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$ (aquamarine line) and $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ (dark-blue line) as a function of $T/NL$; evaluating them in correspondence of $h_{r}^{\text{RHM}}=1/2$ (the related, critical density value is indicated in light-blu) returns the values $e^{-{h_{r}^{\text{RHM}}}}=1/\sqrt{e}\simeq 0.606$ and $e^{-{2h_{r}^{\text{RHM}}}}=e^{-1}\simeq 0.367$; evaluating them in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ (the related, critical density value is indicated in magenta) returns the values $e^{-{h_{f}^{\text{RHM}}}}=1/L=0.001$ and $e^{-{2h_{f}^{\text{RHM}}}}=1/L^{2}=10^{-6}$.
-The red line indicates the density value corresponding to $h=1$, ensuring that each, expected hyperdegree amounts to 1.
-Right panel: trends of $p_{0}$, as defined by Eq.
-(26), and $P_{0}$, i.e.
-the percentage of pairs of parallel hyperedges.
-As already noticed, the shift on the right of the trend induced by the HCM is probably due to the presence of small fitnesses, decreasing the magnitude of the set of probability coefficients.
-The dense (sparse) regime is recovered for large (small) values of $T/NL$.
-Each dot represents an average taken over an ensemble of $1000$ configurations (explicitly sampled from either the RHM or the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
-The yellow line indicates the density value in correspondence of which the minimum, expected hyperdegree amounts to 1.
-The resolution threshold.
-The value of the resolution threshold has been determined by inspecting the asymptotic behaviour of $\langle d_{\alpha\beta}\rangle$: evaluating $P(d_{\alpha\beta}=0)\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $e^{-1}\simeq 0.367$.
-Upon considering that $p_{0}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating it in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $1/\sqrt{e}\simeq 0.606$: in words, the resolution threshold ensures the likelihood of observing any, two parallel hyperedges to be almost half of the probability of each, single hyperedge to be empty, hence suggesting that getting rid of parallel hyperedges is ‘easier’ than getting rid of empty ones.
-Crossing the thresholds.
-As Fig.
-3 confirms, the probability for any, two hyperedges to be parallel decreases faster than the probability of each, single hyperedge to be empty: in words, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges, when empty hyperedges are no longer observed, parallel hyperedges are no longer observed, too.
-More formally, $h_{r}^{\text{RHM}}=1/2<h_{f}^{\text{RHM}}=\ln L$: hence, progressively rising the density of 1s in the incidence matrix leads us crossing the resolution threshold before crossing the filling threshold.
-In more quantitative terms, evaluating $\langle N_{0}\rangle\xrightarrow[]{N\to+\infty}Le^{-h}$ in correspondence of $h_{r}^{\text{RHM}}=1/2$ returns the value $L/\sqrt{e}\simeq 606$ which is larger than $\langle d_{\alpha\beta}\rangle\xrightarrow[]{N\to+\infty}2h$ evaluated in correspondence of $h_{r}^{\text{RHM}}=1/2$ returning the value 1 by definition.
-IV.1.2 2. The Hypergraph Configuration Model
-Let us, now, focus on the trends induced by our heterogeneous benchmarks.
-Following the procedure described in the previous sections, both the fitnesses of nodes and those of hyperedges were drawn from a Pareto distribution [^2] Other fat-tailed distributions were considered: qualitatively, results do not change.
-with $\alpha=2$.
-The right panel of Fig.
-2 depicts the trend of $p_{1}$, as defined by Eq.
-(27).
-The right panel of Fig.
-3 depicts the trend of $p_{0}$, as defined by Eq.
-(26), and the arithmetic mean $P_{0}\equiv 2\sum_{\alpha<\beta}P(d_{\alpha\beta}=0)/L(L-1)$ with $P(d_{\alpha\beta}=0)=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})$, i.e.
-the average of the value $P(d_{\alpha\beta}=0)$ over all pairs of hyperedges - a choice dictated by the will of accounting for the heterogeneity of different pairs of hyperedges.
-Both quantities have been plotted as a function of $\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$.
+23 and 4 represents an average taken over an ensemble of $10^{3}$ configurations explicitly sampled from the RHM and is accompanied by the corresponding $95\%$ confidence interval, calculated via the bootstrap method [41].
+The filling threshold
+Panel a of Fig.
+2 depicts the (analytical) trend of $\langle N_{\emptyset}\rangle/L=p_{\emptyset}=(1-p)^{N}$ (solid line): its agreement with the numerical estimations (dots) confirms the correctness of our formula.
+Although the value of the filling threshold has been determined by inspecting the asymptotic behaviour of $\langle N_{\emptyset}\rangle$, the quantity showing the neatest transition from the sparse to the dense regime is the probability of observing at least one empty hyperedge
+${\color[rgb]{0,0,0}P(N_{\emptyset}>0)=1-P(N_{\emptyset}=0)=1-(1-p_{\emptyset})^{L}=1-[1-(1-p)^{N}]^{L}},$ (52)
+where we have exploited the fact that $P(N_{\emptyset}>0)$ is nothing but the complementary of the probability that no hyperedge is empty.
+Since $p_{\emptyset}\xrightarrow[]{N\to+\infty}e^{-h}$, evaluating such an expression in correspondence of $h_{f}^{\text{RHM}}=\ln L\simeq 6.907$ returns the value $1/L=10^{-3}$ (see panel a of Fig.
+2).
+As a consequence, the result $P(N_{\emptyset}>0)\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$ is numerically recovered (see panel b of Fig.
+2).
+In words, although the filling threshold ensures that each, single hyperedge is empty with an overall small probability, the likelihood of observing at least one, empty hyperedge is still large (i.e.
+2/3$\simeq$ 2/3)≃ 2 / 3 ): the steepness of the trend of $P(N_{\emptyset}>0)$, however, suggests it to quickly vanish as the density of $1$s in the incidence matrix crosses the value $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$.
+Figure 2: Impact of empty hyperedges on the RHM ensemble.
+More in detail, the trends of $\langle N_{\emptyset}\rangle/L=p_{\emptyset}=(1-p)^{N}\xrightarrow[]{N\to+\infty}e^{-h}$, i.e.
+the probability for the generic hyperedge to be emtpy is represented in panel a and the one of $P(N_{\emptyset}>0)=1-[1-(1-p)^{N}]^{L}\xrightarrow[]{N\to+\infty}1-(1-e^{-h})^{L}$, i.e.
+the probability of observing at least one empty hyperedge is represented in panel b.
+Evaluating them in correspondence of $p_{f}^{\text{RHM}}=h_{f}^{\text{RHM}}/N=\ln L/N\simeq 0.023$ (vertical line) returns, respectively, the values $1/L=10^{-3}$ and $1-(1-1/L)^{L}\simeq 0.6323$.
+The dense (sparse) regime is recovered for large (small) values of $p$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
+The multiple resolution threshold
+Panel a of Fig.
+3 depicts the trend of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=[1-2p(1-p)]^{N}=p_{\mathbin{/\!/}}$ (solid line): again, its agreement with the numerical estimations (dots) confirms the correctness of our formula.
+Evaluating $p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$ in correspondence of $h_{m}^{\text{RHM}}\lesssim\ln L\simeq 6.907$ returns the value $1/L^{2}=10^{-6}$ (see panel a of Fig.
+3): in words, the aforementioned, critical value causes the likelihood of observing any two parallel hyperedges to be almost the square of the probability for each single hyperedge to be empty (i.e.
+$1/L$, see the comments under Eq.
+17).
+As the overlaps between pairs of hyperedges cannot be treated as i.i.d.
+random variables, evaluating the probability of observing at least one pair of parallel hyperedges forces us to proceed in a purely numerical fashion.
+Calculating $P(N_{\mathbin{/\!/}}>0)$ in correspondence of $p_{m}^{\text{RHM}}=h_{m}^{\text{RHM}}/N\lesssim\ln L/N\simeq 0.023$ returns the value $0.6$ (see the panel b of Fig.
+3).
+Carrying out such an estimation in the aforementioned regime is, however, instructive as it leads to the expression $P(N_{\mathbin{/\!/}}>0)=1-(1-p_{\mathbin{/\!/}})^{L(L-1)/2}=1-\{1-[1-2p(1-p)]^{N}\}^{L(L-1)/2}\xrightarrow[]{N\to+\infty}1-[1-e^{-2h}]^{L(L-1)/2}$ that, evaluated in correspondence of $h_{m}^{\text{RHM}}\lesssim\ln L\simeq 6.907$, returns the value $P(N_{\mathbin{/\!/}}>0)=1-(1-1/L^{2})^{L(L-1)/2}\simeq 0.393$ whose difference with $0.6$, obtained numerically, lets us fully appreciate the role played by correlations.
+Figure 3: Impact of parallel hyperedges on the RHM ensemble.
+More in detail, the trends of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=[1-2p(1-p)]^{N}=p_{\mathbin{/\!/}}\xrightarrow[]{N\to+\infty}e^{-2h}$, i.e.
+the probability for the generic pair of hyperedges to be parallel and $P(N_{\mathbin{/\!/}}>0)$, i.e.
+the probability of observing at least one pair of parallel hyperedges, as functions of $p$, are represented, respectively, in panels a and b.
+Evaluating them in correspondence of $p_{m}^{\text{RHM}}=h_{m}^{\text{RHM}}/N\lesssim\ln L/N\simeq 0.023$ (vertical line) returns, respectively, the values $1/L^{2}=10^{-6}$ and $\simeq 0.6$.
+The dense (sparse) regime is recovered for large (small) values of $p$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
+The percolation threshold
+Let us, now, focus on the projection of our hypergraph onto the layer of hyperedges.
+The generic hyperedge is isolated either because is not ‘connected’ with any node or because is a singleton (i.e.
+it is ‘connected’ with a node with which no other hyperedge is ‘connected’): in symbols,
+${\color[rgb]{0,0,0}p_{0}\equiv\{1-p[1-(1-p)^{L-1}]\}^{N}=[(1-p)+p(1-p)^{L-1}]^{N}}.$ (53)
+In order to evaluate $p_{0}$ in correspondence of the percolation threshold, let us consider that $L=s^{2}N$, with $s^{2}=10/3$; one, then, finds
+${\color[rgb]{0,0,0}\lim_{N\to+\infty}p_{0}=\lim_{N\to+\infty}\left[\left(1-\frac{1}{Ns}\right)+\frac{1}{Ns}\left(1-\frac{1}{Ns}\right)^{s^{2}N-1}\right]^{N}=e^{-\frac{1-e^{-s}}{s}}},$ (54)
+whose numerical value amounts to $0.631$ (see panel a of Fig.
+4): in words, the value $p_{0}(p=p_{p}^{\text{RHM}})\lesssim 2/3$ implies that the expected number of isolated hyperedges in the projection $\langle N_{0}\rangle=Lp_{0}$ tends to $32L/$ 3 as $p$ tends to $p_{p}^{\text{RHM}}$.
+Figure 4: Impact of isolated hyperedges on the RHM ensemble.
+More in detail, trends of $\langle N_{0}\rangle/L=p_{0}$, i.e.
+the probability for the generic hyperedge to be isolated in the projection, $P(N_{0}>0)$, i.e.
+the probability of observing at least one, isolated hyperedge in the projection and $|\text{LCC}|/N$, i.e.
+the percentage of nodes belonging to the largest connected component (LCC), are represented as functions of $p$, respectively in panels a, b and c.
+Evaluating the first two in correspondence of $p_{p}^{\text{RHM}}=h_{p}^{\text{RHM}}/N=1/\sqrt{NL}\simeq 0.002$ (vertical line) returns, respectively, the values $e^{(e^{-s}-1)/s}\simeq 0.631$ and $\simeq 1$.
+The dense (sparse) regime is recovered for large (small) values of $p$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the RHM) and is accompanied by the corresponding $95\%$ confidence interval.
+Pairs of hyperedges cannot be treated as independent.
+Let us, in fact, consider the bipartite representation of a hypergraph: as adjacent pairs of hyperedges (say $\alpha$, $\beta$ and $\beta$, $\gamma$) may share some neighbours (on the opposite layer), the number of common neighbours of $\alpha$ and $\beta$ will, in general, covariate with the number of common neighbours of $\beta$ and $\gamma$; therefore, evaluating the probability of observing at least one, isolated hyperedge in the projection forces us to proceed in a purely numerical fashion.
+Calculating $P(N_{0}>0)$ in correspondence of $p_{p}^{\text{RHM}}=h_{p}^{\text{RHM}}/N=1/\sqrt{NL}\simeq 0.002$ practically returns $1$ (see panel b of Fig.
+4).
+From the perspective of a hypergraph connectedness, the percolation threshold is ‘less strict’ than the filling threshold, allowing for a larger number of disconnected nodes ($2/3$ of the total versus $1$).
+As before, estimating the percolation threshold in the regime where the pairs of hyperedges behave as i.i.d.
+Binomial random variables is instructive.
+In this case, projecting a bipartite networks onto the layer of hyperedges amounts to connect any two of them with probability $1-(1-p^{2})^{N}$ i.e.
+the complementary of the probability $(1-p^{2})^{N}$ of not sharing any node.
+The number of isolated nodes, thus, obeys the relationship $N_{0}\sim\text{Bin}(L,p_{0})$, with
+${\color[rgb]{0,0,0}p_{0}\equiv[(1-p^{2})^{N}]^{L}}$ (55)
+being the probability for the generic hyperedge to be isolated: in words, such an expression returns the probability for the generic hyperedge to not share any node - with probability $(1-p^{2})^{N}$ with any other hyperedge - with a probability amounting to the previous one raised to the power of $L$.
+As a consequence, evaluating $p_{0}$ in correspondence of the percolation threshold returns a value tending to $1/3$, a result further implying that the expected number of isolated hyperedges in the projection $\langle N_{0}\rangle=Lp_{0}$ tends to the value $L/3$ both letting us fully appreciate the role played by correlations.
+Within such a context, the probability of observing at least one isolated hyperedge in the projection satisfies the chain of relationships $P(N_{0}>0)=1-P(N_{0}=0)=1-(1-p_{0})^{L}=1-[1-(1-p^{2})^{NL}]^{L}\xrightarrow[]{N\to+\infty}1-(1-e^{-1})^{L}$: as the last expression quickly converges to 1 for large values of $L$, the same qualitative behaviour observed before is thus recovered.
+Finally, one may wonder which kind of mesoscale structure is identified by the percolation threshold: the answer is provided by panel c of Fig.
+4, showing the appearance of a large connected component - as also pointed out in [12], the presence of a large connected component is inspected by projecting our hypergraph onto the layer of nodes, whose connectedness is ensured by requiring the connectedness of hyperedges.
+The role of thresholds in a hypergraph evolution
+Let us, now, make a couple of observations.
+The first one concerns the result according to which $h_{m}^{\text{RHM}}\lesssim h_{f}^{\text{RHM}}$: such a relationship suggests that, while filling a large portion of the incidence matrix is not required to observe a limited amount of parallel hyperedges (see panel b of Fig.
+3), when empty hyperedges are no longer observed, parallel hyperedges are no longer observed as well.
+The second one concerns the result according to which either $h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$ or $h_{f}^{\text{RHM}}\geq h_{p}^{\text{RHM}}$.
+By progressively rising the parameter $p$, two, different thresholds are, thus, met: if $h_{f}^{\text{RHM}}\leq h_{p}^{\text{RHM}}$, the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$ is met before the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{NL}$, i.e.
+hyperedges are filled before they start sharing nodes - as a consequence, singletons appear; if $h_{f}^{\text{RHM}}\geq h_{p}^{\text{RHM}}$, the percolation threshold $p_{p}^{\text{RHM}}=1/\sqrt{NL}$ is met before the filling threshold $p_{f}^{\text{RHM}}=\ln L/N$, i.e.
+hyperedges start sharing nodes before they are filled - as a consequence, no singleton appears before the filling threshold is crossed.
+Notice that, for simple graphs, $k_{p}^{\text{RGM}}=1\leq k_{c}^{\text{RGM}}=\ln N$, i.e.
+by progressively rising the parameter $q$, the percolation threshold is always met before the connectivity threshold.
+The Hypergraph Configuration Model
+In order to carry out the numerical simulations in the case of the HCM, we have followed the procedure described in the previous sections and drawn both the fitnesses of nodes and those of hyperedges from a Pareto distribution with $\alpha=2$ other fat-tailed distributions were considered: qualitatively, results do not change.
+Each quantity has been plotted as a function of $\rho(z)=\langle T\rangle/NL\in[10^{-6},1]$ where $\langle T\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}zf_{i}g_{\alpha}/(1+zf_{i}g_{\alpha})$ varies with $z$.
 The dense (sparse) regime is recovered for large (small) values of $z$.
 Each dot of Figs.
-2 and 3 represents an average taken over an ensemble of $1000$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval.
-As in the case of the RHM, the agreement between the analytical trends (solid lines) and the numerical estimations (dots) confirms the correctness of our formulas.
-Comparing the analytical estimation of (the trend of) the percentage of pairs of parallel hyperedges with its sample counterpart reveals the former to provide a quite accurate estimation of the latter.
-Besides, we explicitly notice that the trends characterising the HCM are sistematically shifted on the right: this is probably due to the presence of small fitnesses that, at least in case of Fig.
-2, increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $p_{1}$ vanish.
-Such a result seems to be confirmed by the trends depicted in Fig.
-3.
-Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of 1s in the incidence matrix guaranteeing that the minimum, expected hyperdegree amounts to 1: as evident from Fig.
-3, such a threshold still lies in the right tail of the trends induced by the HCM - although its, precise, numerical value depends on those of the fitnesses.
-IV.2 B. Solving the HCM on real-world hypergraphs
+56 and 7 represents an average taken over an ensemble of $10^{3}$ configurations explicitly sampled from the HCM and is accompanied by the corresponding $95\%$ confidence interval.
+The filling threshold
+Panel a of Fig.
+5 depicts the (analytical) trend of $\langle N_{\emptyset}\rangle/L=\sum_{\alpha=1}^{L}p_{\emptyset}^{\alpha}/L=\overline{p_{\emptyset}}$ (solid line): as in the case of the RHM, its agreement with the numerical estimations (dots) confirms the correctness of our formula.
+Deriving an explicit expression for the filling threshold in the case of the HCM is a rather difficult task; still, we can proceed in a purely numerical fashion and individuate the value of the density of $1$s in the incidence matrix guaranteeing that the expected number of empty hyperedges (divided by $L$) amounts to $1$ (divided by $L$): although its, precise, numerical value depends on the values of the fitnesses, such a threshold still lies in the right tail of the trend induced by the HCM and reads $p_{f}^{\text{HCM}}\simeq 0.032$ (see panel a of Fig.
+5).
+Figure 5: Impact of empty hyperedges on the HCM ensemble.
+More in detail, trends of $\langle N_{\emptyset}\rangle/L=\overline{p_{\emptyset}}$, i.e.
+the probability for the generic hyperedge to be empty and $P(N_{\emptyset}>0)\simeq 1-\prod_{\alpha=1}^{L}(1-e^{-h_{\alpha}})$, i.e.
+the probability of observing at least one empty hyperedge, are represented as functions of the connectance $\rho$, respectively in panels a and b.
+Evaluating the latter in correspondence of the filling threshold, reading $p_{f}^{\text{HCM}}\simeq 0.032$ (vertical line), returns the value $\simeq 0.642$.
+The dense (sparse) regime is recovered for large (small) values of $z$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
+Even in this case, the quantity showing the neatest transition from the sparse to the dense regime is the probability of observing at least one empty hyperedge
+${\color[rgb]{0,0,0}P(N_{\emptyset}>0)=1-P(N_{\emptyset}=0)=1-\prod_{\alpha=1}^{L}(1-p_{\emptyset}^{\alpha})=1-\prod_{\alpha=1}^{L}\left[1-\prod_{i=1}^{N}(1-p_{i\alpha})\right]}$ (56)
+that, in the sparse regime, can be approximated as $P(N_{\emptyset}>0)\simeq 1-\prod_{\alpha=1}^{L}(1-e^{-h_{\alpha}})$: as panel b of Fig.
+5 shows, evaluating $P(N_{\emptyset}>0)$ in correspondence of the filling threshold returns $0.642$.
+Finally, let us explicitly notice that the value of the filling threshold is shifted on the right with respect to its homogeneous counterpart, an evidence probably due to the presence of small fitnesses that increase the probability of observing at least one, empty hyperedge, hence requiring a larger value of $z$ to let $P(N_{\emptyset}>0)$ vanish.
+Figure 6: Impact of parallel hyperedges on the HCM ensemble.
+More in detail, trends of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=\overline{p_{\mathbin{/\!/}}}$, i.e.
+the probability for the generic pair of hyperedges to be parallel and $P(N_{\mathbin{/\!/}}>0)$, i.e.
+the probability of observing at least one pair of parallel hyperedges, are represented as functions of the connectance $\rho$, respectively in panels a and b.
+Evaluating the latter in correspondence of the multiple resolution threshold, reading $p_{m}^{\text{HCM}}\simeq 0.031$ (vertical line), returns the $0.493$.
+The dense (sparse) regime is recovered for large (small) values of $z$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
+The multiple resolution threshold
+Panel a of Fig.
+6 depicts the (analytical) trend of $2\langle N_{\mathbin{/\!/}}\rangle/L(L-1)=2\sum_{\alpha=1}^{L}\sum_{\begin{subarray}{c}\beta=1\\ \beta>\alpha\end{subarray}}^{L}p_{\mathbin{/\!/}}^{\alpha\beta}/L(L-1)=\overline{p_{\mathbin{/\!/}}}$ (solid line): as in the case of the RHM, its agreement with the numerical estimations (dots) confirms the correctness of our formula.
+Adopting the strategy described in the previous paragraph, i.e.
+that of requiring that the expected number of parallel hyperedges (divided by $L(L-1)/2$) amounts to $1$ (divided by $L(L-1)/2$), we found that $p_{m}^{\text{HCM}}\simeq 0.031$ (see panel a of Fig.
+6).
+Numerically clculating $P(N_{\mathbin{/\!/}}>0)$ in correspondence of the multiple resolution threshold returns the value $0.493$ (see panel b of Fig.
+6).
+Let us notice that the value of the multiple resolution threshold no longer coincides with the value of the filling threshold, although it is still shifted on the right with respect to its homogeneous counterpart.
+Figure 7: Impact of isolated hyperedges on the HCM ensemble.
+More in detail, trends of $\langle N_{0}\rangle/L=\overline{p_{0}}$, i.e.
+the probability for the generic hyperedge to be isolated in the projection, $P(N_{0}>0)$, i.e.
+the probability of observing at least one, isolated hyperedge in the projection, and $|\text{LCC}|/N$, i.e.
+the percentage of nodes belonging to the largest connected component (LCC), are repesented as functions of the connectance $\rho$, respectively in panels a, b and c.
+Evaluating the first two in correspondence of $p_{p}^{\text{HCM}}\simeq 0.001$ (vertical line), returns the values $\simeq 0.766$ and $\simeq 1$.
+The dense (sparse) regime is recovered for large (small) values of $p$.
+Each dot represents an average taken over an ensemble of $10^{3}$ configurations (explicitly sampled from the HCM) and is accompanied by the corresponding $95\%$ confidence interval.
+The percolation threshold
+The probability for the generic hyperedge to be isolated in the projection, now, reads
+${\color[rgb]{0,0,0}p_{0}^{\alpha}\equiv\prod_{i=1}^{N}\left\{1-p_{i\alpha}\left[1-\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right]\right\}=\prod_{i=1}^{N}\left\{(1-p_{i\alpha})+p_{i\alpha}\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right\}}$ (57)
+that, in the sparse regime, can be approximated as $p_{0}^{\alpha}\simeq\prod_{i=1}^{N}[(1-p_{i\alpha})+p_{i\alpha}e^{-k_{i}}]$.
+Adopting the strategy described in the previous paragraphs, i.e.
+that of requiring that the expected value of the total number of nodes shared by any hyperedge with any other hyperedge amounts to $1$ in symbols, $\overline{\langle\sigma\rangle}=\sum_{\alpha=1}^{L}\langle\sigma_{\alpha}\rangle/L=1$ we found that $p_{p}^{\text{HCM}}\simeq 0.001$: since $\langle N_{0}\rangle=\sum_{\alpha=1}^{L}p_{0}^{\alpha}$, evaluating $\langle N_{0}\rangle/L=\sum_{\alpha=1}^{L}p_{0}^{\alpha}/L=\overline{p_{0}}$ and $P(N_{0}>0)$ in correspondence of the percolation threshold returns respectively the values $0.766$ (see panel a of Fig.
+7) and $\simeq 1$ (see panel b of Fig.
+7).
+For what concerns the hypergraph connectedness, the same conclusion drawn in the case of the RHM holds true, as the percolation threshold allows for a larger number of disconnected nodes ($3/4$ of the total versus $1$).
+Analogously, the mesoscale structure individuated by the percolation threshold consists of a large connected component constituted (see panel c of Fig. 7).
+Solving the HCM on real-world hypergraphs
 In order to test our benchmarks on real-world configurations, we have focused on a number of data sets taken from Austin R.
-Benson’s website [^3] https://www.cs.cornell.edu/~arb/data/, i.e.
+Benson’s website (https://www.cs.cornell.edu/~arb/data/), i.e.
 the contact-primary-school, the email-Enron and the NDC-classes ones.
-To have an explicit expression for the expected value of any quantity of interest, let us comment on the goodness of the approximation
-$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha.$ (34)
-As Fig.
-7 in Appendix A shows, it is is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
-IV.3 C. ‘Hypergraph to graph’ projection
+Although the parameters defining the HCM must be numerically determined by solving the system of equations induced by the likelihood maximisation, when the system under analysis is sparse they can be approximated as described in A of the Appendix.
+Such an approximation leads to the expression
+$p_{i\alpha}\simeq x_{i}y_{\alpha}=\frac{k_{i}^{*}h_{\alpha}^{*}}{T^{*}},\>\forall\>i,\alpha$ (58)
+that, as Fig.
+11 in the Appendix shows, is quite accurate for each data set considered here - in fact, one can safely assume that $x_{i}\simeq k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
+‘Hypergraph to graph’ projection
 The canonical formalism that we have adopted leads to factorisable distributions, i.e.
 distributions that can be written as a product of pair-wise probability distributions; this allows the expectation of several quantities of interest to be evaluated analytically.
-To this aim, let us consider the matrix introduced in Battiston et al.
-(2020), reading
+Let us start by considering the matrix, introduced in [4], reading
 $\mathbf{W}=\mathbf{I}\cdot\mathbf{I}^{T}-\mathbf{K}$ (59)
-with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above
-$w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}-\delta_{ij}k_{i}$ (36)
-counts the number of hyperedges nodes $i$ and $j$ are connected by - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$.
-The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
-$\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}\rangle\langle I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha};$ (37)
-the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 4) can be computed as
-$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (38)
-whose expected value reads
-$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (39)
-As further confirmed by Fig.
-8 in Appendix A, the approximation provided by Eq.
-(34) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$.
-Interestingly, as Fig.
-5a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
-Let us, now, extend the concept of assortativity to hypergraphs. To this aim, we consider the quantity named average incident hyperedges degree, defined as
-$k_{i}^{nn}=\sum_{\alpha=1}^{L}\frac{I_{i\alpha}h_{\alpha}}{k_{i}}=\frac{\sigma_{i}+k_{i}}{k_{i}}=\frac{\sigma_{i}}{k_{i}}+1\simeq\frac{\sigma_{i}}{k_{i}}$ (40)
-and representing the arithmetic mean of the degrees of the hyperedges including node $i$. An analytical approximation of its expected value can be provided as well:
-$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}$
-$=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (41)
+with $\mathbf{K}$ being the diagonal matrix whose $i$-th entry reads $k_{i}$; according to the definition above, it induces a projection of a hypergraph onto a weighted graph, whose generic entry
 Figure 8: Graphical representation of the ‘hypergraph to graph’ projection.
@@
 the number of nodes it shares an hyperedge with or, equivalently, the degree on the projection) reads $\{\kappa_{i}\}=(5,4,5,4,5,3)$.
-IV.4 D. Disparity ratio and degree in the projection
+$w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}-\delta_{ij}k_{i}$ (60)
+returns the number of hyperedges both $i$ and $j$ belong to - more explicitly, $w_{ij}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$, $i\neq j$ and $w_{ii}=\sum_{\alpha=1}^{L}I_{i\alpha}I_{i\alpha}-k_{i}=\sum_{\alpha=1}^{L}I_{i\alpha}-k_{i}=k_{i}-k_{i}=0$.
+In other words, $\mathbf{W}$ represents the object most closely resembling a traditional adjacency matrix.
+The null models discussed so far can be employed to calculate $\langle w_{ij}\rangle$, $i\neq j$ that, in a perfectly general fashion, reads
+$\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}\rangle\langle I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha};$ (61)
+as we said, the total number of hyperedges shared by node $i$ with any other node in the hypergraph (in a sense, its ‘strength’ - see also Fig. 8) can be computed as
+$\sigma_{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}w_{ij}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\sum_{\alpha=1}^{L}I_{i\alpha}I_{j\alpha}$ (62)
+whose expected value reads
+$\langle\sigma_{i}\rangle$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\langle w_{ij}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}[\langle h_{\alpha}\rangle-p_{i\alpha}].$ (63)
+As further confirmed by Fig.
+11 in the Appendix, the approximation provided by Eq.
+(58) allows us to pose $\langle\sigma_{i}\rangle_{\text{HCM}}\simeq k_{i}^{*}\sum_{\alpha=1}^{L}(h_{\alpha}^{*})^{2}/T^{*}$, $\forall\>i$.
+Interestingly, as Fig.
+9a shows, the HCM overestimates the extent to which any two nodes of the email-Enron data set overlap: in words, such a real-world hypergraph is more compartmentalised than expected.
+Let us, now, extend the concept of assortativity to hypergraphs. To this aim, we consider the quantity named average incident hyperedges degree, defined as
+$k_{i}^{nn}=\sum_{\alpha=1}^{L}\frac{I_{i\alpha}h_{\alpha}}{k_{i}}=\frac{\sigma_{i}+k_{i}}{k_{i}}=\frac{\sigma_{i}}{k_{i}}+1\simeq\frac{\sigma_{i}}{k_{i}}$ (64)
+and representing the arithmetic mean of the degrees of the hyperedges including node $i$. An analytical approximation of its expected value can be provided as well:
+$\langle k_{i}^{nn}\rangle$ $\simeq\sum_{\alpha=1}^{L}\frac{p_{i\alpha}[\langle h_{\alpha}\rangle+1-p_{i\alpha}]}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle+\langle k_{i}\rangle}{\langle k_{i}\rangle}=\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}+1\simeq\frac{\langle\sigma_{i}\rangle}{\langle k_{i}\rangle}.$ (65)
+Disparity ratio and degree in the projection
 More information about the patterns shaping real-world hypergraphs can be obtained upon defining the ratio $f_{ij}=w_{ij}/\sigma_{i}$, $i\neq j$ that induces the quantity
-$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (42)
-known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig.
+$Y_{i}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}f_{ij}^{2}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{w_{ij}^{2}}{\sigma_{i}^{2}},$ (66)
+known as disparity ratio and quantifying the (un)evenness of the distribution of the weights constituting the strength of node $i$ over the $\kappa_{i}=\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\Theta[w_{ij}]\equiv\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}a_{ij}$ links characterising its connectivity - since $a_{ij}=1$ if nodes $i$ and $j$ share, at least, one hyperedge, $\kappa_{i}$ is the degree of node $i$ in the projection of the hypergraph (see also Fig.
 8 and Fig.
-5b).
+9b).
 Since, under the RHM, $w_{ij}\sim\text{Bin}(L,p^{2})$, we find that $\langle a_{ij}\rangle=1-(1-p^{2})^{L}$, i.e.
@@
 An analytical approximation of the expected value of the disparity ratio of node $i$ can be provided as well:
-$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ (j\neq i)\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (45)
-Figure 5: Scatter plots between the empirical values $\{\sigma_{i}\}$, $\{\kappa_{i}\}$, $\{Y_{i}\}$, $\{\text{CEC}_{i}\}$ and the expected values $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$, $\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$, $\langle\{Y_{i}\rangle_{\text{HCM}}\}$, $\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$.
+$\langle Y_{i}\rangle\simeq\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}.$ (69)
+Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed.
+In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig.
+13 in the Appendix).
+For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM.
+In any case, as Fig.
+9c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected.
+A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}\left[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})\right]$: as Fig.
+9b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected.
+This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs [42, 43].
+Figure 9: Scatter plots between the empirical and the expected values for the email-Enron dataset.
+More in detail, $\{\sigma_{i}\}$ vs.
+$\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$ (panel a), $\{\kappa_{i}\}$ vs.
+$\{\langle\kappa_{i}\rangle_{\text{HCM}}\}$ (panel b), $\{Y_{i}\}$ vs.
+$\langle\{Y_{i}\rangle_{\text{HCM}}\}$ (panel c), and $\{\text{CEC}_{i}\}$ vs.
+$\{\langle\text{CEC}_{i}\rangle_{\text{HCM}}\}$ (panel d).
 The HCM overestimates the extent to which any, two nodes overlap, as well as the CEC; the disparity ratio, instead, is underestimated by it.
 These results can be understood by considering that the HCM just constrains the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed.
-Contrarily to what has been previously observed, the expected value of the disparity ratio cannot be always safely decomposed as a ratio of expected values - not even if the ‘full’ HCM is employed.
-In fact, while this approximation works relatively well for the contact-primary-school data set, it does not for the email-Enron and the NDC-classes ones (see also Fig.
-10 in Appendix C).
-For this reason, the expected value of the disparity ratio has been evaluated by explicitly sampling the ensemble of incidence matrices induced by the ‘full’ HCM.
-In any case, as Fig.
-5c shows, such a null model underestimates the disparity ratio characterising each node of the email-Enron data set: in words, the empirical overlap between any two nodes is (much) less evenly ‘distributed’ than expected.
-A similar conclusion can be drawn by considering $\langle\kappa_{i}\rangle=\sum_{i=1}^{N}[1-\prod_{\alpha=1}^{L}(1-p_{i\alpha}p_{j\alpha})]$: as Fig.
-5b shows, the degree of the nodes in the projection tend to be significantly smaller than expected, meaning that hyperedges concentrate on fewer edges than expected.
-This observation is in line with recent works showing the encapsulation and ‘simpliciality’ of real-world hypergraphs LaRock and Lambiotte (2023); Landry et al.
-(2024).
-IV.5 F. Eigenvector centrality
+Eigenvector centrality
 Centrality measures for hypergraphs have been defined as well.
-An example is provided by the clique motif eigenvector centrality (CEC), defined in Benson (2019) (see also Appendix D): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$.
+An example is provided by the clique motif eigenvector centrality (CEC), defined in [44] (see also Section C of the Appendix): $\text{CEC}_{i}$ corresponds to the $i$-th entry of the Perron-Frobenius eigenvector of $\mathbf{W}$.
 As Fig.
-5d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
-Figure 6: Top panels: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes.
-Bottom panels: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges.
+9d shows, the HCM underestimates the CEC as well: such a result can be understood by considering that the HCM constrains only the degree sequences, hence inducing an ensemble where connections are ‘distributed’ more evenly than observed, an evidence letting the nodes overlap more, thus causing the entries of $\langle\mathbf{W}\rangle$ to be overall larger and less dissimilar, as well as those of its Perron-Frobenius eigenvector.
+Confusion matrix
+Let us, now, consider the set of indices constituting the so-called confusion matrix (see also Section D of the Appendix).
+They are intended to quantify the capability of a given network model in reproducing microscopic properties such as the position of $1$s and 00s by explicitly comparing their empirical location with the one expected under the chosen model.
+They are named true positive rate (TPR), i.e.
+the percentage of $1$s correctly recovered by a given method, whose expected value reads
+$\langle\text{TPR}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{T};$ (70)
+specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method, whose expected value reads
+$\langle\text{SPC}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{(1-I_{i\alpha})(1-p_{i\alpha})}{NL-T};$ (71)
+positive predictive value (PPV), i.e. the percentage of $1$s correctly recovered by a given method with respect to the total number of $1$s predicted by it, whose expected value reads
+$\langle\text{PPV}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{\langle T\rangle};$ (72)
+accuracy (ACC), measuring the overall performance of a given method in correctly placing both $1$s and 00s, whose expected value reads
+$\langle\text{ACC}\rangle=\frac{\langle\text{TP}\rangle+\langle\text{TN}\rangle}{NL}.$ (73)
+Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 2 in the Appendix); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of $1$s - and, as a consequence, the density of 00s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 3 in Section E of the Appendix).
+Figure 10: Validated vs.
+non-validated ’hypergraph to network’ projections of empirical datasets.
+Panels a, b and c: projections of the contact-primary-school, email-Enron and NDC-classes data sets onto the layer of nodes.
+Panels d, e and f: validated counterparts of the aforementioned projections: any two nodes are linked if they share a significantly large number of hyperedges.
 Communities have been detected by running the Louvain algorithm.
-IV.6 G. Confusion matrix
-Let us, now, consider the set of properties composing the so-called confusion matrix (see also Appendix E).
-They are known as the true positive rate (TPR), i.e.
-the percentage of 1s correctly recovered by a given method and whose expected value reads
-$\langle\text{TPR}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{T};$ (46)
-the specificity (SPC), i.e. the percentage of 0s correctly recovered by a given method and whose expected value reads
-$\langle\text{SPC}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{(1-I_{i\alpha})(1-p_{i\alpha})}{NL-T};$ (47)
-the positive predictive value (PPV), i.e. the percentage of 1s correctly recovered by a given method with respect to the total number of 1s predicted by it and whose expected value reads
-$\langle\text{PPV}\rangle=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{\langle T\rangle};$ (48)
-and the accuracy (ACC), measuring the overall performance of a given method in correctly placing both 1s and 0s and whose expected value reads
-$\langle\text{ACC}\rangle=\frac{\langle\text{TP}\rangle+\langle\text{TN}\rangle}{NL}.$ (49)
-Results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter ones makes it difficult to reproduce the TPR and the PPV (see also Table 3 in Appendix E); on the other hand, the capability of the HCM (both in its ‘full’ and approximated version) to reproduce the density of 1s - and, as a consequence, the density of 0s - ensures the SPC to be recovered quite precisely, in turn ensuring the overall accuracy of the model to be large (for an overall evaluation of the performance of the HCM in reproducing real-world hypergraphs, see also Table 4 in Appendix F).
-IV.7 H. Community detection
+Community detection
 Communities are commonly understood as densely connected groups of nodes.
 Representing an hypergraph via its incidence matrix allows this statement to be made more precise from a statistical perspective: in fact, the null models discussed so far can be employed to test if any two nodes share a significantly large number of hyperedges - hence can be clustered together, should this be the case.
-To this aim, we can adapt the recipe proposed in Saracco et al. (2017) to project bipartite networks and summarised in the following. One, first, computes
+In other words, it is possible to devise a ‘validation procedure’ that filters the projection described by the matrix $\mathbf{W}$ by removing the entries that do not satisfy the requirement above.
+To this aim, we can adapt the recipe proposed in [24] to project bipartite networks and summarised in the following. One, first, computes
 […]
@@
 Third, one links the (pairs of) nodes whose related p-value is smaller than the aforementioned threshold.
-Figure 6 shows the result returned by the Louvain algorithm, run on the validated projections: as noticed elsewhere Saracco et al.
-(2017); Pratelli et al.
-(2024); Guarino et al.
-(2024), the detection of mesoscale structures is enhanced if carried out on filtered topologies.
-V V. Conclusions
+Fig.
+10 shows the partitions returned by the Louvain algorithm run on the validated projections: as noticed elsewhere [24, 46, 47], the detection of mesoscale structures is enhanced if carried out on filtered topologies.
+Conclusions
 Our paper contributes to current research on hypergraphs by extending the constrained entropy-maximisation framework to incidence matrices, i.e.
 their simplest, tabular representation.
-Differently from the currently-available techniques Chodrow (2020), our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
+Differently from the currently-available techniques [9], our methodology has the advantage of being analytically tractable, scalable and versatile enough to be straightforwardly extensible to directed and/or weighted hypergraphs.
 Beside leading to results whose relevance is mostly theoretical (i.e.
@@
 for detecting patterns that are not imputable to purely random effects.
-Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Appendix F).
+Specifically, our results suggest that real-world hypergraphs are characterised by a degree of self-organisation that is absolutely non-trivial (see also Section E of the Appendix).
 This is even more surprising when considering that our results are obtained under a benchmark such as the HCM, i.e.
-a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes de Arruda et al.
-(2021); Iacopini et al.
-(2022); St-Onge et al.
-(2022) - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
-VI VI. Acknowledgements
+a null model constraining both the degree and the hyperdegree sequences: since it overestimates the extent to which any two nodes overlap - a result whose relevance becomes evident as soon as one considers the effects that higher-order structures have on spreading and cooperation processes [48, 49, 50] - our future efforts will be directed towards the analysis of benchmarks constraining non-linear quantities such as the co-occurrences between nodes and/or hyperedges.
+[…]
+Acknowledgements
 RL acknowledges support from the EPSRC grants EP/V013068/1, EP/V03474X/1 and EP/Y028872/1.
@@
 FS acknowledges support from the project ‘CODE - Coupling Opinion Dynamics with Epidemics’, funded under PNRR Mission 4 ‘Education and Research’ - Component C2 - Investment 1.1 - Next Generation EU ‘Fund for National Research Program and Projects of Significant National Interest’ PRIN 2022 PNRR, grant code P2022AKRZ9.
+Declarations
+Competing interests
+The authors declare no competing interests.
+Data availability
+All the data used for the analysis are freely available at https://www.cs.cornell.edu/~arb/data/.
+Code availability
+The authors will provide the code used for the analysis upon request.
+Author contribution
+F.S., G.P., R.L.
+and T.S.
+conceived the study.
+F.S.
+performed the analyses and T.S.
+wrote the first version of the manuscript.
+F.S., G.P., R.L.
+and T.S.
+discussed the results, revised the draft and approved the final version of the manuscript.
+Appendix A The Hypergraph Configuration Model
+A.1 Approximating the HCM
+Upon considering that the generic probability coefficient induced by the HCM can be Taylor-expanded as
+$p_{i\alpha}^{\text{HCM}}=\frac{x_{i}y_{\alpha}}{1+x_{i}y_{\alpha}}=x_{i}y_{\alpha}-(x_{i}y_{\alpha})^{2}+(x_{i}y_{\alpha})^{3}\dots$ (76)
+one may wonder at which order the expansion can be safely truncated.
+In case hypergraphs are sparse, i.e.
+$T/NL\ll 1$, and in absence of hubs, it turns out that just considering the first order is enough to obtain predictions as accurate as those achievable under the ‘full’ HCM: in other words, one can simply put $p_{i\alpha}^{\text{FOA}}\simeq x_{i}y_{\alpha}$.
+Upon doing so, the system of equations defining the HCM simplifies to
+$k_{i}^{*}$ $=\sum_{\alpha=1}^{L}x_{i}y_{\alpha},\>\forall\>i$ (77)
+$h_{\alpha}^{*}$ $=\sum_{i=1}^{N}x_{i}y_{\alpha},\>\forall\>\alpha$ (78)
+expressions leading us to find $x_{i}=k_{i}^{*}/\sqrt{T^{*}}$, $\forall\>i$ and $y_{\alpha}=h_{\alpha}^{*}/\sqrt{T^{*}}$, $\forall\>\alpha$.
+As a consequence, $p_{i\alpha}^{\text{FOA}}=x_{i}y_{\alpha}=k_{i}^{*}h_{\alpha}^{*}/T^{*}$ a position that is commonly named Chung-Lu approximation (CLA).
+Fig.
+11 shows the goodness of the latter in approximating $\langle\sigma_{i}\rangle_{\text{HCM}}$ for the contact-primary-school, the email-Enron and the congress-bills data sets, i.e.
+the accuracy of the following chain of equalities
+$\langle\sigma_{i}\rangle_{\text{HCM}}$ $=\sum_{\alpha=1}^{L}p_{i\alpha}^{\text{HCM}}[h_{\alpha}-p_{i\alpha}^{\text{HCM}}]$
+[…]
+[…]
+$\simeq k_{i}^{*}\sum_{\alpha=1}^{L}\frac{(h_{\alpha}^{*})^{2}}{T^{*}},\>\forall\>i.$ (79)
+Figure 11: Effectiveness of Chung-Lu approximation.
+Panels a, d and g: numerical values of the node-specific Lagrange multipliers obtained by solving the ‘full’ HCM scattered versus the ones derived by adopting the CLA.
+Panels b, e and h: numerical values of the hyperedge-specific Lagrange multipliers obtained by solving the ‘full’ HCM scattered versus the ones derived by adopting the CLA.
+Panels c, f and i: expected values $\{\langle\sigma_{i}\rangle_{\text{HCM}}\}$ scattered versus the expected values $\{\langle\sigma_{i}\rangle_{\text{CLA}}\}$.
+Table 1: Summary of the constraints defining the canonical version of each benchmark considered here.
+While the total number of $1$s is preserved by each model, the degrees are separately preserved by the two, partial HCMs and jointly preserved only by the HCM.
+Stars indicate the empirical values of the constraints.
+$\langle T\rangle$ $\langle k_{i}\rangle$ $\langle h_{\alpha}\rangle$
+Random Hypergraph Model $T^{*}$ $T^{*}/N$ $T^{*}/L$
+Partial Configuration Model - nodes layer $T^{*}$ $k_{i}^{*}$ $T^{*}/L$
+Partial Configuration Model - hyperedges layer $T^{*}$ $T^{*}/N$ $h_{\alpha}^{*}$
+Hypergraph Configuration Model $T^{*}$ $k_{i}^{*}$ $h_{\alpha}^{*}$
+A.2 Partial Configuration Models
+The HCM reduces to a ‘partial’ Configuration Model in case one of the two degree sequences is left unconstrained.
+Partial Configuration Model - nodes layer.
+Let us start by solely constraining the degree sequence $\{k_{i}\}_{i=1}^{N}$, by posing $\beta_{\alpha}=0$, $\forall\>\alpha$ (or, equivalently, $y_{\alpha}=1$, $\forall\>\alpha$).
+Our canonical probability distribution becomes
+$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}\frac{x_{i}^{I_{i\alpha}}}{1+x_{i}}=\prod_{i=1}^{N}p_{i}^{k_{i}}(1-p_{i})^{L-k_{i}}$ (80)
+where $e^{-\alpha_{i}}\equiv x_{i}$ and $p_{i}\equiv x_{i}/(1+x_{i})$, $\forall\>i$.
+The entries of the incidence matrix are, now, independent random variables obeying the Bernoulli distributions reading
+$I_{i\alpha}\sim\text{Ber}(p_{i}),\>\forall\>i,\alpha;$ (81)
+in words, the entries along the same row obey the same distribution while the ones along the same column obey different distributions.
+As a consequence, the $i$-th node degree (a sum of i.i.d.
+Bernoulli random variables), obeys the Binomial distribution
+$k_{i}\sim\text{Bin}(L,p_{i})$ (82)
+while the $\alpha$-th hyperedge degree (a sum of independent random variables that obey different Bernoulli distributions) obeys the Poisson-Binomial distribution
+$h_{\alpha}\sim\text{PoissBin}(N,p_{1}\dots p_{N}).$ (83)
+Interestingly, while the degrees of the nodes obey different Binomial distributions, the degrees of the hyperedges obey the same Poisson-Binomial distribution.
+The resolution of the likelihood maximisation problem leads us to find the values $p_{i}=k_{i}^{*}/L$, $\forall\>i$ which, in turn, ensure that $\langle k_{i}\rangle_{\text{HPCM}_{\text{nodes}}}=\sum_{\alpha=1}^{L}p_{i}=k_{i}^{*}$, $\forall\>i$ and that $\langle T\rangle_{\text{HPCM}_{\text{nodes}}}=T^{*}$; instead, $\langle h_{\alpha}\rangle_{\text{HPCM}_{\text{nodes}}}$ will, in general, differ from $h_{\alpha}^{*}$ in fact, $\langle h_{\alpha}\rangle_{\text{HPCM}_{\text{nodes}}}=T^{*}/L$, $\forall\>\alpha$.
+From a microcanonical perspective, the number of configurations satisfying the requirement that the degrees of the nodes match their empirical values amounts to
+$\Omega_{\text{HPCM}_{\text{nodes}}}=\prod_{i=1}^{N}\binom{L}{k_{i}^{*}};$ (84)
+reshuffling the $1$s along each row of the incidence matrix separately ensures the degrees of the nodes to be preserved while destroying any other correlation.
+Partial Configuration Model - hyperedges layer.
+Analogously, the canonical probability distribution describing the case in which only the hyperdegree sequence $\{h_{\alpha}\}_{\alpha=1}^{L}$ is constrained reads
+$P(\mathbf{I})=\prod_{i=1}^{N}\prod_{\alpha=1}^{L}\frac{y_{\alpha}^{I_{i\alpha}}}{1+y_{\alpha}}=\prod_{\alpha=1}^{L}p_{\alpha}^{h_{\alpha}}(1-p_{\alpha})^{N-h_{\alpha}}$ (85)
+where $e^{-\beta_{\alpha}}\equiv y_{\alpha}$ and $p_{\alpha}\equiv y_{\alpha}/(1+y_{\alpha})$, $\forall\>\alpha$.
+As for the previous null model, the entries of the incidence matrix are independent random variables obeying different Bernoulli distributions, i.e.
+$I_{i\alpha}\sim\text{Ber}(p_{\alpha}),\>\forall\>i,\alpha;$ (86)
+in words, the entries along the same column obey the same distribution while the ones along the same row obey different distributions.
+As a consequence, the $i$-th node degree (a sum of independent random variables that obey different Bernoulli distributions) obeys the Poisson-Binomial distribution
+$k_{i}\sim\text{PoissBin}(L,p_{1}\dots p_{L})$ (87)
+while the $\alpha$-th hyperedge degree (a sum of i.i.d. Bernoulli random variables) obeys the Binomial distribution
+$h_{\alpha}\sim\text{Bin}(N,p_{\alpha}).$ (88)
+Interestingly, while the degrees of the nodes obey the same Poisson-Binomial distribution, the degrees of the hyperedges obey different Binomial distributions.
+The resolution of the likelihood maximisation problem leads us to find the values $p_{\alpha}=h_{\alpha}^{*}/N$, $\forall\>\alpha$ which, in turn, ensure that $\langle h_{\alpha}\rangle_{\text{HPCM}_{\text{hyperedges}}}=\sum_{i=1}^{N}p_{\alpha}=h_{\alpha}^{*}$, $\forall\>\alpha$ and that $\langle T\rangle_{\text{HPCM}_{\text{hyperedges}}}=T^{*}$; instead, $\langle k_{i}\rangle_{\text{HPCM}_{\text{hyperedges}}}$ will, in general, differ from $k_{i}^{*}$ in fact, $\langle k_{i}\rangle_{\text{HPCM}_{\text{hyperedges}}}=\sum_{\alpha=1}^{L}p_{\alpha}=T^{*}/N$, $\forall\>i$.
+From a microcanonical perspective, the number of configurations satisfying the requirement that the degrees of the hyperedges match their empirical values amounts to
+$\Omega_{\text{HPCM}_{\text{hyperedges}}}=\prod_{\alpha=1}^{L}\binom{N}{h_{\alpha}^{*}};$ (89)
+reshuffling the $1$s along each column of the incidence matrix separately ensures the degrees of the hyperedges to be preserved, while destroying any other correlation.
+[…]
+A.3 Asymptotic results
+As noticed in the main text, the asymptotic behaviour of the HCM depends on the whole set of coefficients $\{p_{i\alpha}\}_{i,\alpha}$: closed form expressions can be, thus, obtained only after having specified the functional dependence of each of them on $N$.
+General conclusions can be drawn more easily for the two, ‘partial’ Configuration Models.
+[…]
+$p_{\emptyset}^{\alpha}$ $=\prod_{i=1}^{N}(1-p_{i\alpha})=\left(1-\frac{h_{\alpha}}{N}\right)^{N}\rightarrow e^{-h_{\alpha}},$ (90)
+$p_{\mathbin{/\!/}}^{\alpha\beta}$ $=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})\simeq\left(1-\frac{h_{\alpha}+h_{\beta}}{N}\right)^{N}\rightarrow e^{-(h_{\alpha}+h_{\beta})},$ (91)
+[…]
+$p_{0}^{\alpha}=\prod_{i=1}^{N}\left\{1-p_{i\alpha}+p_{i\alpha}\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right\}\simeq\left[1-\frac{h_{\alpha}}{N}(1-e^{-T/N})\right]^{N}\rightarrow e^{-h_{\alpha}(1-e^{-T/N})},$ (92)
+whose compatibility with the prediction returned by the ‘full’ HCM has been inspected only numerically (see Fig. 12).
+[…]
+$p_{\emptyset}^{\alpha}$ $=\prod_{i=1}^{N}(1-p_{i\alpha})=\prod_{i=1}^{N}\left(1-\frac{k_{i}}{L}\right)\simeq\prod_{i=1}^{N}e^{-\frac{k_{i}}{L}}=e^{-\frac{\sum_{i=1}^{N}k_{i}}{L}}=e^{-\frac{T}{L}}=e^{-h},$ (93)
+$p_{\mathbin{/\!/}}^{\alpha\beta}$ $=\prod_{i=1}^{N}(1-q_{i}^{\alpha\beta})=\prod_{i=1}^{N}\left(1-\frac{2k_{i}}{L}\right)\simeq\prod_{i=1}^{N}e^{-\frac{2k_{i}}{L}}=e^{-\frac{2\sum_{i=1}^{N}k_{i}}{L}}=e^{-\frac{2T}{L}}=e^{-2h},$ (94)
+[…]
+$p_{0}^{\alpha}=\prod_{i=1}^{N}\left\{1-p_{i\alpha}+p_{i\alpha}\prod_{\begin{subarray}{c}\beta=1\\ \beta\neq\alpha\end{subarray}}^{L}(1-p_{i\beta})\right\}\simeq\prod_{i=1}^{N}\left[1-\frac{k_{i}}{L}\left(1-e^{-k_{i}}\right)\right]\simeq e^{-\frac{\sum_{i=1}^{N}k_{i}(1-e^{-k_{i}})}{L}},$ (95)
+whose compatibility with the prediction returned by the RHM has been inspected only numerically (see Fig. 12).
+Appendix B Expected value of topological quantities
+Let us, now, provide the explicit expression of the expected value of some topological properties of interest.
+Let us start with the entries of the matrix $\mathbf{W}$, for which the following results hold true
+$\langle w_{ij}\rangle$ $=\sum_{\alpha=1}^{L}\langle I_{i\alpha}I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}\langle I_{i\alpha}\rangle\langle I_{j\alpha}\rangle=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha},$ (96)
+$\text{Var}[w_{ij}]$ $=\sum_{\alpha=1}^{L}\text{Var}[I_{i\alpha}I_{j\alpha}]+2\sum_{\beta=1}^{L}\sum_{\begin{subarray}{c}\gamma=1\\ \gamma>\beta\end{subarray}}^{L}\text{Cov}[I_{i\beta}I_{j\beta},I_{i\gamma}I_{j\gamma}]=\sum_{\alpha=1}^{L}p_{i\alpha}p_{j\alpha}(1-p_{i\alpha}p_{j\alpha}).$ (97)
+Therefore,
+$\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\langle w_{ij}\rangle^{2}$ $=\sum_{\alpha=1}^{L}p_{i\alpha}^{2}\left[\sum_{j=1}^{N}p_{j\alpha}^{2}-p_{i\alpha}^{2}\right]+2\sum_{\beta=1}^{L}\sum_{\begin{subarray}{c}\gamma=1\\ \gamma>\beta\end{subarray}}^{L}p_{i\beta}p_{i\gamma}\left[\sum_{j=1}^{N}p_{j\beta}p_{j\gamma}-p_{i\beta}p_{i\gamma}\right],$ (98)
+$\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\text{Var}[w_{ij}]$ $=\sum_{\alpha=1}^{L}p_{i\alpha}(\langle h_{\alpha}\rangle-p_{i\alpha})-\sum_{\alpha=1}^{L}p_{i\alpha}^{2}\left[\sum_{j=1}^{N}p_{j\alpha}^{2}-p_{i\alpha}^{2}\right].$ (99)
+Figure 13: The limits of analytical estimates of the disparity ratio under HCM.
+Panels a, d and g: scatter plots between the analytical and the sample estimates of the numerator of the disparity ratio, under the ‘full’ HCM.
+Panels b, e and h: scatter plots between the analytical and the sample estimates of the denominator of the disparity ratio, under the ‘full’ HCM.
+Panels c, f and i: scatter plots between the expected values $\{\langle Y_{i}\rangle_{\text{HCM}}\}$, computed as in Eq.
+(102), and the ones obtained by explicitly sampling the ensemble induced by the ‘full’ HCM.
+As analytical and sampled values of numerators and denominators nicely agree, the result concerning $\{\langle Y_{i}\rangle\}$ must be imputed to the correlation between the two, disregarded by Eq.
+(102).
+Let us, now, consider the strength sequence of matrix $\mathbf{W}$, for which the following results hold true
+$\langle\sigma_{i}\rangle$ $=\sum_{\alpha=1}^{L}p_{i\alpha}(\langle h_{\alpha}\rangle-p_{i\alpha}),$ (100)
+$\text{Var}[\sigma_{i}]$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\text{Var}[w_{ij}]+2\sum_{\begin{subarray}{c}j<k\\ j,k\neq i\end{subarray}}\text{Cov}[w_{ij},w_{ik}]$
+$=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\text{Var}[w_{ij}]+2\sum_{\begin{subarray}{c}j<k\\ j,k\neq i\end{subarray}}[\langle w_{ij}w_{ik}\rangle-\langle w_{ij}\rangle\langle w_{ik}\rangle].$ (101)
+As a consequence, the expected value of the disparity ratio reads
+$\langle Y_{i}\rangle$ $\simeq\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{\langle w_{ij}^{2}\rangle}{\langle\sigma_{i}^{2}\rangle}=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\frac{\langle w_{ij}\rangle^{2}+\text{Var}[w_{ij}]}{\langle\sigma_{i}\rangle^{2}+\text{Var}[\sigma_{i}]}$ (102)
+where the $\simeq$ symbol is understood to approximate the expected value of a ratio as the ratio of expected values. This approximation just represents the first order of the Taylor expansion
+$\text{E}\left[\frac{X}{Y}\right]\simeq\frac{\text{E}[X]}{\text{E}[Y]}-\frac{\text{Cov}[X,Y]}{\text{E}[Y]^{2}}+\frac{\text{E}[X]}{\text{E}[Y]^{3}}\text{Var}[X];$ (103)
+due to the correlation between numerator and denominator, the second and the third terms of Eq.
+(103) may not be negligible: as Fig.
+13 shows, while Eq.
+(102) holds true for the contact-primary-school data set, it does not for the others.
+Appendix C Expected value of the clique motif eigenvector centrality
+Figure 14: Comparison between the ensemble average and the analytical estimate of CEC.
+Scatter plot between the ensemble average (of each component) of the Perron-Frobenius eigenvector of a given, empirical matrix $\mathbf{W}^{*}$ and the numerical value (of each component) of the Perron-Frobenius eigenvector of the corresponding, expected matrix $\langle\mathbf{W}\rangle_{\text{HCM}}$ for contact-primary-school, email-Enron and NDC-classes datasets (respectively, panels a, b and c).
+The extremely good agreement suggests a faster way to calculate the CEC that avoids to explicitly sample the ensemble induced by a null model.
+The clique motif eigenvector centrality (CEC) has been proposed in [44]: its $i$-th entry, $\text{CEC}_{i}$, is nothing else that the corresponding entry of the Perron-Frobenius eigenvector of $\mathbf{W}$.
+In order to evaluate the expected value of the CEC, we have explicitly sampled the ensemble of incidence matrices induced by the HCM; then, we have calculated the matrix $\tilde{\mathbf{W}}$ induced by each $\tilde{\mathbf{I}}$, according to Eq.
+(59) of the main text.
+Afterwards, we have calculated the Perron-Frobenius eigenvector of each ‘projected’ matrix and taken their average in an entry-wise fashion.
+Remarkably, we found that the aforementioned ensemble average basically coincides with the Perron-Frobenius eigenvector of the ensemble average of $\mathbf{W}$ itself, i.e.
+$\langle\mathbf{W}\rangle_{\text{HCM}}$, as Fig.
+14 shows.
+Appendix D Expected value of the entries of the confusion matrix
+Table 2: Confusion matrix and density of $1$s for the three real-world hypergraphs considered in the present paper.
+[…]
+contact-primary-school $3.57\times 10^{-8}$ 0.991 $3.57\times 10^{-8}$ 0.983 $8.66\times 10^{-3}$
+email-Enron $8.15\times 10^{-8}$ 0.982 $8.15\times 10^{-8}$ 0.966 $1.72\times 10^{-2}$
+NDC-classes $2.40\times 10^{-10}$ 0.997 $2.40\times 10^{-10}$ 0.995 $2.71\times 10^{-3}$
+Before defining the entries of the confusion matrix, let us call $\tilde{\mathbf{I}}$ the generic member of the ensemble of incidence matrices induced by one of the benchmarks considered in the present paper: the expected value of any quantity which is a function of $\tilde{\mathbf{I}}$ is readily calculated by averaging it over the ensemble itself.
+The true positive rate (TPR) is defined as the percentage of $1$s correctly recovered by a given model and reads
+$\text{TPR}=\frac{\text{TP}}{T}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}\tilde{I}_{i\alpha}}{T}\Longrightarrow\langle\text{TPR}\rangle=\frac{\langle\text{TP}\rangle}{T}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{T};$ (104)
+the specificity (SPC) is defined as the percentage of 00s correctly recovered by a given model and reads
+SPC $=\frac{\text{TN}}{NL-T}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{(1-I_{i\alpha})(1-\tilde{I}_{i\alpha})}{NL-T}\Longrightarrow$
+[…]
+the positive predictive value (PPV) is defined as the percentage of $1$s correctly recovered by a given model with respect to the total number of predicted of $1$s and reads
+$\text{PPV}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}\tilde{I}_{i\alpha}}{\tilde{T}}\Longrightarrow\langle\text{PPV}\rangle=\frac{\langle\text{TP}\rangle}{\langle T\rangle}=\sum_{i=1}^{N}\sum_{\alpha=1}^{L}\frac{I_{i\alpha}p_{i\alpha}}{\langle T\rangle};$ (106)
+lastly, the accuracy (ACC) measures the overall performance of a given reconstruction method in correctly placing both $1$s and 00s and reads
+$\text{ACC}=\frac{\text{TP}+\text{TN}}{NL}\Longrightarrow\langle\text{ACC}\rangle=\frac{\langle\text{TP}\rangle+\langle\text{TN}\rangle}{NL}.$ (107)
+As stressed in the main text, results on the confusion matrix of a number of real-world hypergraphs reveal that the large sparsity of the latter makes it difficult to reproduce the TPR and the PPV (see Table 2); on the other hand, the capability of the HCM to reproduce the density of $1$s makes it capable of recovering the density of 00s as well, thus ensuring the overall accuracy of the model to be quite large.
+Appendix E Comparing the RHM and the HCM performance
+Table 3: The $R^{2}$ index allows us to compare the performance of the RHM with that of the HCM in reproducing the patterns of our real-world hypergraphs.
+While both models perform worse than the baseline model, the HCM is, overall, more accurate than the RHM in reproducing all quantities we have considered, the only exception being $\sigma$: therefore, just constraining the total number of $1$s of a given incidence matrix is, in general, not enough to achieve an accurate description of the hypergraph it represents - although constraining the degree and the hyperdegree sequences may be not enough as well.
+Data set $R^{2}_{\text{HCM}}(\sigma)$ $R^{2}_{\text{HCM}}(Y)$ $R^{2}_{\text{HCM}}(\text{CEC})$ $R^{2}_{\text{RHM}}(\sigma)$ $R^{2}_{\text{RHM}}(Y)$ $R^{2}_{\text{RHM}}(\text{CEC})$
+contact-primary-school -2.9697 -1.7248 -0.1489 -2.7349 -1.8189 -0.8478
+email-Enron -0.1344 -0.2591 -0.0519 -0.0131 -0.3462 -0.1705
+NDC-classes -0.0044 -0.1107 -0.0482 -0.0008 -0.1222 -0.0183
+Let us, now, inspect the performance of the RHM and that of the HCM in reproducing the patterns of our real-world hypergraphs.
+To this aim, let us consider the coefficient of determination that, in our case, can be re-defined as
+$R^{2}_{\text{null}}(\mathbf{X})=1-\frac{\sum_{i=1}^{N}\left(X_{i}-\langle X_{i}\rangle_{\text{null}}\right)^{2}}{\sum_{i=1}^{N}\left(X_{i}-\bar{X}\right)^{2}};$ (108)
+it measures the goodness of a null model in reproducing empirical data when compared to a baseline model whose only prediction reads $\bar{X}$, i.e.
+the arithmetic mean of the empirical data themselves.
+While such a model is characterised by $R^{2}=0$, a model that is capable of matching the observed values exactly is characterised by $R^{2}=1$; models returning predictions whose discrepancies are larger than the discrepancies accompanying the predictions of the baseline model are characterised by $R^{2}<0$.
+As Table 3 shows, the HCM is often more accurate than the RHM in reproducing the quantities we have considered.
+More precisely, the RHM is found to steadily perform worse than it when the CEC is considered and to perform quite similar to it when the $Y$ is considered; the opposite tendency is observed when the $\sigma$ is considered.
+Overall, these results confirm that just constraining the total number of $1$s of a given incidence matrix is not enough to achieve an accurate description of the corresponding hypergraph but also suggest that constraining the degree and the hyperdegree sequences may be not enough as well - as we have already noticed when commenting the highly non-trivial degree of self-organisation of the real-world hypergraphs (whose trends are) depicted in Fig.
+9 of the main text.